### v1

# Anomalous and linear holographic hard wall models for light unflavored mesons

Abstract
In this work we consider anomalous and linear holographic hard wall (HW) models for light unflavored mesons inspired by the AdS/CFT correspondence. The anomalous dimensions depend on the logarithm of the spin $S$ of the meson state and come from a semiclassical analysis of gauge/string duality. The anomalous HW model produces very good masses and good Regge trajectories for mesons compared with PDG data. Inspired by this anomalous HW model we also propose a phenomenological modification of the dimension of the boundary operators such that the model produces asymptotic linear Regge trajectories.
I Introduction
The AdS/CFT correspondence Aharony:1999ti establishes relations between Yang-Mills (YM) and higher dimensional gravitational theories. One of the amazing properties of the correspondence is that if YM theories are strongly coupled, the gravitational one is weak coupled allowing for the solution of a myriad of phenomena from the quark-gluon plasma to condensed matter (for reviews, see e.g., Casalderrey-Solana:2011dxg ; Zaanen:2015oix ). In particular, hadronic spectra were described by different proposals as the holographic hard wall (HW) and the soft wall (SW) models Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us . The HW model Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct is well known for producing reasonable hadronic masses but non-linear (parabolic) Regge trajectories. On the other side, the SW model Karch:2006pv ; Colangelo:2008us also give good masses with the advantage of generating linear Regge trajectories.
In this work, we are going to consider the inclusion of anomalous dimensions for the boundary operators describing light unflavored mesons with spins $S\geq 1$. These anomalous dimensions depend on the logarithm of the spin $S$ of the mesonic state inspired by a semiclasical analysis of gauge/string dualities Gubser:2002tv and was used recently to discuss the case of even spins glueballs Costa-Silva:2023vuu . We find very good masses and Regge trajectories for the studied mesons when compared with PDG data PDG:2022pth , better than the ones from the HW and SW models. Furthermore, we discuss a phenomenological modification of the dimension of the boundary operators representing the mesonic states in order to obtain asymptotic linear Regge trajectories, similar to the case of glueballs. Actually, the linearization of the mesonic states require an additional parameter to obtain the best fit to PDG data. For a recent account on non-holographic meson Regge trajectories, see e.g. Chen:2021kfw .
This work is organized as follows. In Section II, we review the HW model for light unflavored mesons and in Section III we discuss the introduction of anomalous dimensions in the HW model and obtain the spectra for the mesonic families starting with the states of the $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. Then, in Section IV we modify the dimension of the mesonic boundary operators such that we find asymptotic linear Regge trajectories for the mesons already discussed in the previous section. Finally, in Section V we present our conclusions.
II Original HW model for mesons
In this work we are interested in studying the spectra of light unflavored mesons, and their Regge trajectories using anomalous and linear holographic HW models. These mesons can be classified in four families according to their behavior under $\mathcal{CP}$ transformations (even or odd), isospin $I=0$ and $1$, and spins $S\geq 1$. These four families imply four distinct Regge trajectories.
In order to describe such objects using the holographic HW model, we need to study a massive vector field in $AdS_{5}$. We start with the action
$S=-\frac{1}{2}\int d^{5}x\sqrt{-g}\bigg{[}\frac{1}{2}g^{pm}g^{qn}F_{mn}F_{pq}+M_{5}^{2}g^{pm}A_{p}A_{m}\bigg{]},$ (1)
where $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$ is the stress tensor of the Proca field with mass $M_{5}$ in the five-dimensional AdS space. Moreover, $g$ is the determinant of the metric of $AdS_{5}$,
$ds^{2}=g_{mn}dx^{m}dx^{n}=\frac{R^{2}}{z^{2}}(dz^{2}+\eta_{\mu\nu}dx^{\mu}dx^{\nu}),$ (2)
with $R$ being the $AdS$ radius, $z$ the holographic coordinate and $\eta_{\mu\nu}$ is the Minkowski four-dimensional boundary metric. The equation of motion that follows from the action, Eq. (1), is given by
$\partial_{z}\bigg{[}\bigg{(}\frac{1}{z}\bigg{)}F_{zn}\eta^{nq}\bigg{]}+\partial_{\mu}\bigg{[}\bigg{(}\frac{1}{z}\bigg{)}\eta^{m\mu}F_{mn}\eta^{nq}\bigg{]}-\frac{(M_{5}R)^{2}}{z^{3}}A_{n}\eta^{nq}=0.$ (3)
We use a plane wave solution as an ansatz for the five-dimensional vector field $A_{m}(z,x^{\mu})$ with four-momentum $q_{\mu}$, and polarization four-vector $\epsilon_{\rho}$,
$A_{\rho}(z,x^{\mu})=\epsilon_{\rho}v(z)e^{iq_{\mu}x^{\mu}},$ (4)
$A_{z}=0,$ (5)
with $\partial_{\mu}A^{\mu}=0$. Defining $v(z)=zf(z)$, one obtains
$z^{2}\frac{d^{2}\psi(z)}{dz^{2}}+z\frac{d\psi(z)}{dz}-q^{2}z^{2}\psi(z)-[1+(M_{5}R)^{2}]f(z)=0.$ (6)
This is Bessel equation, whose solutions are linear combination of Bessel ($J_{\nu}$) and Neumann ($Y_{\nu})$ functions of order $\nu=\sqrt{1+(M_{5}R)^{2}}$. Since we are interested only in regular solutions in the bulk, we will consider just the Bessel solution
$f(z)\sim J_{\nu}(mz),$ (7)
where $m$ is an irrelevant constant in the conformal field theory defined on the Minkowski boundary of the $AdS_{5}$ space, due to scale invariance, but has a fundamental role for the HW model with the introduction of the cutoff in the AdS space (see Eq. (15) below).
On the other hand, the mesons with spin $S$ with $S=1,2,3,...$, that we are interested in are described by operators in the boundary Minkowski space as Gubser:2002tv  [^1] It is also possible to alternatively write the mesons with spin $S$ as deTeramond:2005su ${\cal O}^{\mu\,\mu_{1}...\mu_{S}}_{2+S}=\;\bar{\psi}\gamma^{\mu}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (8) such that $\Delta=2+S.$ (9) However, in this case, the results for the AHW and LAHW are not so good as the ones discussed in Sections III and IV, with $\Delta=3+S$.
${\cal O}^{\mu_{1}...\mu_{S}}_{3+S}=\;\bar{\psi}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (10)
where $\psi$ and $\bar{\psi}$ represent quark and antiquark fields while $D_{\mu_{i}}$ with $i=1,\dots,S$ are symmetrized covariant derivatives. Then, the canonical dimensions of these operators are given by
$\Delta=\Delta_{\rm can.}=3+S.$ (11)
The link between the operators $\cal O$ and string excitations at $AdS_{5}$ can be constructed from its conformal dimension. From the AdS/CFT correspondence, the vector field in $AdS_{5}$ that we are describing is a $1$-form, so its holographic mass $M_{5}$ is related to the conformal dimension $\Delta$ of boundary operators $\cal O$, is given by
$(M_{5}R)^{2}=(\Delta-1)(\Delta-3).$ (12)
Using this relation in the index $\nu=\sqrt{1+(M_{5}R)^{2}}$ of the Bessel function, one sees that
$\nu=\Delta-2.$ (13)
Furthermore, combining Eqs. (13) and (11), one finds that the Bessel functions describing the mesons represented by the operators $\cal O$, Eq. (10), have order
$\nu=S+1.$ (14)
The HW model consists in the holographic relation between fields in AdS space with an IR cutoff that breaks the conformal symmetry of the boundary theory. That means a hard cutoff in the holographic $z$ coordinate, so we are working in an AdS slice, given by $0\leq z\leq z_{\rm max}$, where
$z_{\rm max}=\frac{1}{\Lambda_{\rm QCD}},$ (15)
with $\Lambda_{\rm QCD}$ a convenient energy scale related to be fixed later. Then the solution of the vector bulk field, Eq. (7), assumes the form
$f(z)\sim J_{\nu}(m_{\nu,k}z),$ (16)
where $m_{\nu,k}$ is identified with the mass of the boundary operator $\cal O$ that lives in the four-dimensional Minkowski space, describing mesons. Note that $k=1,2,3...$ denote radial excitations, where $k=1$ is the corresponding ground state. In this work we are interested in angular momenta excitations, so we take $k=1$ for all the objects from now on.
Imposing Dirichlet boundary conditions (for $k=1$) in Eq. (16) at $z=z_{\rm max}$, one obtains
$J_{\nu}\bigg{(}\frac{m_{\nu,1}}{\Lambda_{\rm QCD}}\bigg{)}=0.$ (17)
Defining the first zero $\chi_{\nu,1}$ of the Bessel function of order $\nu$, such that $J_{\nu}(\chi_{\nu,1})=0$, and making use of the relation Eq. (14), we can write the mass of the mesonic state with spin $S$ as
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=S+1).$ (18)
In particular, from the state with $S=1$ we fix
$\Lambda_{\rm QCD}=\frac{m_{2,1}}{\chi_{2,1}},$ (19)
taking $m_{2,1}$ as an input from PDG for the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. The corresponding values of $\Lambda_{\rm QCD}$ found here for each mesonic family are shown in Table 2. The meson masses obtained from the HW model, using Eq. (18), are shown in Table 1, together with the SW model predictions with the relative deviations with respect to PDG data PDG:2022pth . In Table 2, we also show the equations for the Regge trajectories found from PDG data for the same mesonic families.
The process described above is known as the holographic HW model and was used in the past to obtain hadronic spectra among other properties. It is a well-known fact that this model implies non-linear Regge trajectories.
In the next section we develop an improvement of the HW model, by considering anomalous dimensions, in order to obtain better masses compared with the PDG ones and better Regge trajectories for these meson families.
Table 1: Masses in MeV for light unflavored mesons classified under even or odd $\mathcal{CP}$ and with isospin $I=0,1$, and spins $S\geq 1$ from PDG PDG:2022pth , from the HW Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct and SW Karch:2006pv ; Colangelo:2008us models, and the percentual deviations $\delta_{\rm HW}$ and $\delta_{\rm SW}$ of the masses from these two models compared with PDG data. The HW masses are determined from Eq. (18), using the values of $\Lambda_{\rm QCD}$ for each family (same $\mathcal{CP}$ parity and Isospin) shown in Table 2, and the masses from the SW model come from the relation $m_{S}^{2}\sim S$ Karch:2006pv . The mass of the state with $J=S=1$ for each family is used as an input for the HW and SW models.
$J^{PC}$ ${CP}$ $I$ Meson PDG HW $\delta_{\rm HW}$ SW $\delta_{\rm SW}$
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 0.25$ $775.26$ $0\%$ $775.26$ $0\%$
$2^{++}$ even 1 $a_{2}(1320)$ $1316.9\pm 0.9$ $963.13$ $25.9\%$ $1096.38$ $16.7\%$
$3^{--}$ even 1 $\rho_{3}(1690)$ $1688.8\pm 2.1$ $1145.52$ $32.2\%$ $1342.79$ $20.5\%$
$4^{++}$ even 1 $a_{4}(1970)$ $1967\pm 16$ $1324.12$ $32.7\%$ $1550.52$ $21.2\%$
$5^{--}$ even 1 $\rho_{5}(2350)$ $2330\pm 35$ $1499.93$ $35.6\%$ $1733.53$ $25.6\%$
$6^{++}$ even 1 $a_{6}(2450)$ $2450\pm 130$ $1673.57$ $31.7\%$ $1898.99$ $22.5\%$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 0.13$ $782.66$ $0\%$ $782.66$ $0\%$
$2^{++}$ even 0 $f_{2}(1270)$ $1275.5\pm 0.8$ $972.31$ $23.8\%$ $1106.85$ $13.2\%$
$3^{--}$ even 0 $\omega_{3}(1670)$ $1667\pm 4$ $1156.44$ $30.6\%$ $1355.61$ $18.7\%$
$4^{++}$ even 0 $f_{4}(2050)$ $2018\pm 11$ $1336.74$ $33.8\%$ $1565.32$ $22.4\%$
$5^{--}$ even 0 $\omega_{5}(2250)$ $2250\pm 70$ $1514.23$ $32.7\%$ $1750.08$ $22.2\%$
$6^{++}$ even 0 $f_{6}(2510)$ $2465\pm 50$ $1689.52$ $31.5\%$ $1917.12$ $22.2\%$
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 3.2$ $1229.5$ $0\%$ $1229.5$ $0\%$
$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1670.6\pm 2.05$ $1527.41$ $8.57\%$ $1738.78$ $4.08\%$
$3^{+-}$ odd 1 $b_{3}(2030)$ $2032\pm 12$ $1816.70$ $10.6\%$ $2129.56$ $4.80\%$
$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2250\pm 15$ $2099.95$ $6.67\%$ $2459.00$ $9.29\%$
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 6$ $1166$ $0\%$ $1166$ $0\%$
$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1617\pm 5$ $1352$ $16.4\%$ $1648.97$ $1.98\%$
$3^{+-}$ odd 0 $h_{3}(2025)$ $2025\pm 20$ $1709$ $15.6\%$ $2019.57$ $0.27\%$
$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2328\pm 38$ $1882$ $19.2\%$ $2332.00$ $0.17\%$
Table 2: The Regge trajectory in GeV units for each light unflavored meson family (same $\mathcal{CP}$ parity and same isospin $I$) from the best fit of PDG data shown in Table 1. We also show the mass of the state with $J=S=1$ for each family, and the corresponding $\Lambda_{\rm QCD}$ value, both in MeV, determined from Eq. (19).
$J^{PC}$ ${CP}$ $I$ Meson PDG (MeV) PDG Regge Trajectory (GeV) $\Lambda_{\rm QCD}$ (MeV)
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 0.25$ $J=(0.88\pm 0.04)M^{2}+0.47\pm 0.15.$ $150.96$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 0.13$ $J=(0.90\pm 0.02)M^{2}+0.47\pm 0.07$ 152.24
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 3.2$ $J=(0.83\pm 0.04)M^{2}-0.30\pm 0.16$ 239.41
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 6$ $J=(0.73\pm 0.02)M^{2}+0.03\pm 0.06$ 227.04
III Anomalous HW model
In an interacting theory, the dimension of an operator is no longer canonical, thanks to renormalization that implies in anomalous contributions ($\Delta_{\rm Anom}$). So, the total dimension of an operator is given by its canonical plus the anomalous dimensions:
$\Delta_{\rm Total}=\Delta_{\rm can.}+\Delta_{\rm Anom}\,.$ (20)
From a well known semi-classical gauge/string duality analysis Gubser:2002tv , it is found that the anomalous dimension increases logarithmically with the spín $S$ of the boundary operators as
$\Delta_{\rm Anom}=\frac{\sqrt{\lambda}}{\pi}\ln{\bigg{(}\frac{S}{\sqrt{\lambda}}\bigg{)}}+\mathcal{O}(S^{0}),$ (21)
where $\lambda$ is ’t Hooft coupling. Inspired by this result, we consider an anomalous dimension given by
$\Delta_{\rm Log}=a\ln(S),$ (22)
where $a$ is a constant that will be determined by best fit for each meson family.
Then, the anomalous HW (AHW) model consists in considering bulk fields in AdS slice, Eq. (15), with total dimension given by Eq. (20) such that the mass associated with boundary operators, now takes the form
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=S+1+\Delta_{\rm Log}),$ (23)
where we used the fact that $\Delta_{\rm can.}=3+S$ for these mesonic states and that the order of the corresponding Bessel function is $\nu=\Delta_{\rm Total}-2$. The value of $\Lambda_{\rm QCD}$ is determined for each meson familiy using as inputs the masses of the states with $S=1$ from PDG, namely $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$. These values of $\Lambda_{\rm QCD}$ coincide with the ones used in the usual HW model, listed in Table 2, once Eq. (22) implies a vanishing anomalous dimension for $S=1$.
In order to obtain the best numerical fit, let us define the percentual deviation:
$q_{i}=\frac{|m_{i}-M_{i}|}{M_{i}},$ (24)
where $m_{i}$ is the mass of a determined particle obtained by our model, from Eq. (23), and $M_{i}$ its mass obtained from PDG, listed in Table 1. Further, we define the function $Q$ as
$Q=\sqrt{\sum_{i}q_{i}^{2}}.$ (25)
So, the best fit method adopted here corresponds to minimize the function $Q$ for each meson family in the AHW model.
We will also consider an uncertainty for the parameter $a$, defined in Eq. (22), which implies in uncertainties for the masses ${m_{i}}$, obtained by error propagation, accordingly to the formula:
$\delta{m_{i}}=\sqrt{\left(\frac{\chi_{\nu_{i},1}}{\chi_{2,1}}\right)^{2}\left(\delta m_{0}\right)^{2}+\left(\frac{m_{0}}{\chi_{2,1}}\right)^{2}\left(\delta\chi_{\nu_{i},1}\right)^{2}},$ (26)
where $\delta m_{0}$ and $\delta\chi_{\nu_{i},1}$ are the uncertainties in the mass of the state with $S=1$ and in the zero of Bessel function that is related to the mass $m_{i}$. The uncertainty $\delta m_{0}$ of the state with $S=1$ for each meson family in our model will be the arithmetic average of the percentual uncertainties of PDG masses $M_{i}$ for that given family. Defined our criterion for best fit and how we compute the uncertainties $\delta{m_{i}}$ for each mass $m_{i}$, we are able to tackle with the AHW model for light unflavored mesons.
III.1 AHW model for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
We start applying the AHW model for mesons with even ${\cal CP}$, isospin $I=1$, and spins $S\geq 1$, taking the mass of the $\rho(770)$ state from PDG as an input, which fixes $\Lambda_{\rm QCD}=150.96$ MeV. We minimize the parameter $Q$, given by Eq. (25), with respect to PDG masses from Table 1, finding the logarithm anomalous dimensions, Eq. (22), to be
$\Delta_{\rm Log}(S)=(2.77\pm 0.12)\ln(S).$ (27)
The masses with its uncertainties obtained within this model, from Eq. (23), with these anomalous dimensions, are presented in Table 3, together with percentual deviations in comparison with PDG data. The value of $\Delta_{\rm Log}(S)$ of the above equation for each spin $S$ is also shown in the last column of this Table. The Regge trajectory obtained by doing linear regression with the obtained masses from this AHW Log model is given by
$J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03,$ (28)
with $\chi^{2}/ndf=4.32/4=1.08$. This trajectory is presented in the upper left panel of Fig. 1 as the blue line, and can be compared with the Regge trajectory of this family from PDG data, shown in Table 2. The red dots are the values in the $J\times M^{2}$ plane of the mesons $\rho(770)$, $a_{2}(1320)$, $\rho_{3}(1690)$, $a_{4}(1970)$, $\rho_{5}(2350)$, and $a_{6}(2450)$ in this AHW Log model, with their uncertainties.
III.2 AHW model for mesons with even ${\cal CP}$, $I=0$, and $S\geq 1$
Here, we discuss the case of mesons with even ${\cal CP}$, isospin $I=0$, and spins $S\geq 1$. Minimizing $Q$, given by Eq. (25), we obtain $a=2.61\pm 0.09$ for the anomalous dimension, Eq. (22), which reads
$\Delta_{\rm log}(S)=(2.61\pm 0.09)\ln(S).$ (29)
Using the mass of the state $\omega(782)$ from PDG as an input, which implies $\Lambda_{\rm QCD}=152.24$ MeV for this case, we calculate the tower of masses with the uncertainties in this AHW Log model and show these results in Table 3. The percentual deviations of these masses in comparison with PDG are also presented, as well as the value of the anomalous dimension $\Delta_{\rm Log}$, for each meson state. The Regge trajectory obtained by doing linear regression with the masses obtained by the AHW Log model is given by
$J=(0.90\pm 0.01)M^{2}+0.48\pm 0.03,$ (30)
with $\chi^{2}/ndf=8.88/4=2.22$. This trajectory is presented in the upper right panel of Fig. 1 as the blue line. The red dots are the values in the $J\times M^{2}$ plane of the mesons with even ${\cal CP}$, isospin $I=0$, and spins $S\geq 1$, in this AHW Log model, with their uncertainties. This trajectory can be compared with the one obtained from PDG for the $\omega(782)$ family, presented in Table 2.
III.3 AHW model for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$
Now we present our results for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$. The minimization of $Q$, Eq. (25), gives $a=0.56\pm 0.03$ for the the anomalous dimension, Eq. (22), which becomes
$\Delta_{\rm Log}(S)=(0.56\pm 0.03)\ln(S).$ (31)
The masses obtained in this AHW Log model are shown in Table 3, with their uncertainties, where we used the mass of the state $b_{1}(1235)$ as an input, so that $\Lambda_{\rm QCD}=239.41$ MeV. The Regge trajectory obtained by linear regression with these masses is given by
$J=(0.78\pm 0.02)M^{2}-0.13\pm 0.08,$ (32)
with $\chi^{2}$/ndf$=27/2=13.5$. This trajectory is represented in the lower left panel of Fig. 1 as the blue line and can be compared with the one obtained from PDG data, shown in Table 2. The red dots are the values obtained from this AHW Log model, with their corresponding uncertainties.
Table 3: Masses in MeV for mesons classified under even or odd $\mathcal{CP}$ with isospin $I=0,1$ and spins $S\geq 1$ from the AHW Log model, determined from Eq. (23), and the percentual deviations $\delta_{\rm Log}$ of the masses from this model compared with PDG data. For each family the mass of the state with $J=S=1$ is taken as an input from PDG. In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq. (22).
$J^{PC}$ $\mathcal{CP}$ $I$ Meson AHW Log $\delta_{\rm Log}$ $\Delta_{\rm Log}$
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$2^{++}$ even 1 $a_{2}(1320)$ $1309.95\pm 22$ $0.53\%$ $1.92$
$3^{--}$ even 1 $\rho_{3}(1690)$ $1681.02\pm 31$ $0.46\%$ $3.04$
$4^{++}$ even 1 $a_{4}(1970)$ $1988.75\pm 37$ $1.11\%$ $3.84$
$5^{--}$ even 1 $\rho_{5}(2350)$ $2262.37\pm 43$ $2.90\%$ $4.46$
$6^{++}$ even 1 $a_{6}(2450)$ $2514.58\pm 47$ $2.64\%$ $4.96$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 7.83$ $0\%$ 00
$2^{++}$ even 0 $f_{2}(1270)$ $1302.58\pm 17$ $2.12\%$ $1.81$
$3^{--}$ even 0 $\omega_{3}(1670)$ $1666.4\pm 23$ $0.04\%$ $2.87$
$4^{++}$ even 0 $f_{4}(2050)$ $1969.63\pm 28$ $2.40\%$ $3.62$
$5^{--}$ even 0 $\omega_{5}(2250)$ $2240.19\pm 32$ $0.44\%$ $4.20$
$6^{++}$ even 0 $f_{6}(2510)$ $2490.23\pm 36$ $1.02\%$ $4.68$
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 5.0$ $0\%$ 00
$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1640.57\pm 8.35$ $1.80\%$ $0.39$
$3^{+-}$ odd 1 $b_{3}(2030)$ $1991.55\pm 11$ $1.99\%$ $0.62$
$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2316.73\pm 13$ $2.97\%$ $0.78$
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 10$ $0\%$ 00
$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1629.93\pm 18$ $0.80\%$ $0.66$
$3^{+-}$ odd 0 $h_{3}(2025)$ $2003.12\pm 25$ $1.08\%$ $1.04$
$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2339.01\pm 30$ $0.47\%$ $1.32$
Figure 1: Regge trajectories (blue lines) for mesons, with masses $m$ expressed in GeV, from the AHW Log model. Red dots are the values obtained from Eq. (23) in the AHW Log model with uncertainties shown in Table 3. The blue lines are obtained by linear regression of these points. Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
III.4 AHW model for mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$
For mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$, we minimize the function $Q$, Eq. (25), finding $a=0.95\pm 0.06$ for the anomalous dimension, Eq. (22), obtaining
$\Delta_{\rm Log}(S)=(0.95\pm 0.06)\ln(S).$ (33)
The masses obtained in this AHW Log model are presented, with uncertainties, in Table 3, where we used the mass value of the state $h_{1}(1170)$ from PDG as an input, such that $\Lambda_{\rm QCD}=227.04$ MeV. The Regge trajectory obtained by doing linear regression with these masses is given by
$J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05,$ (34)
with $\chi^{2}$/ndf$=3.52/2=1.76.$ This trajectory can be compared with the one obtained from PDG data shown in Table 2 and is presented in the lower right panel of Fig. 1 as the blue line. The red dots are the values of the masses obtained within this AHW Log model with their uncertainties.
IV Linear AHW for mesons
Recently, we have shown Costa-Silva:2023vuu that one can obtain asymptotically linear Regge trajectories for glueballs with even spins $S$ from an AHW model, when considering a total dimension $\Delta_{\rm Total}$ proportional to $\sqrt{S}$. In order to find asymptotically linear Regge trajectories for light unflavored mesons with spins $S\geq 1$, we apply a similar procedure here with a small modification in the power of the spins $S$ as
$\Delta_{\rm Lin}=a_{\rm Lin}(S^{\frac{1}{2}+\epsilon}-1)+4,$ (35)
where in the ideal case, $\epsilon\rightarrow 0$. Note that when $S=1$, which are the lowest spin state for each mesonic family considered here, we recover the canonical dimension $\Delta=4$, compatible with Eq. (11).
Here, the parameter $a_{\rm Lin}$ is obtained exactly as in the AHW Log model, by minimizing $Q$ in Eq. (25). The parameter $\epsilon$ that represents a little deviation of the ideal case $\epsilon=0$ is chosen to obtain the best asymptotic linear Regge trajectory for each meson family. The value of $\epsilon$ is the one that minimizes the total deviation of the masses from the straight line representing the Regge trajectory.
The masses that follows from this linear AHW model are determined by
$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=\Delta_{\rm Lin}-2),$ (36)
where $\Delta_{\rm Lin}$ is defined in Eq. (35). Since $\Delta_{\rm Lin}=4$ for $S=1$, the values of $\Lambda_{\rm QCD}$ here in this model, for each meson family, are the same as those in the AHW Log model, as well as in the original HW presented in Table 2.
As in the case of logarithmic anomalous dimension, we will analyze the families of light unflavored mesons with even or odd $\cal CP$, isospins $I=0,1$, and spins $S\geq 1$. The Regge trajectories that we obtain within this model are asymptotically linear for each meson family with the appropriate choice of parameters.
IV.1 Linear AHW for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
For mesons with even ${\cal CP}$, isospin $I=1$, and spin $S\geq 1$, the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=6.87\pm 0.16$ and $\epsilon=0$, so the total dimension reads
$\Delta_{\rm Lin}=(6.87\pm 0.16)\sqrt{S}-2.87\pm 0.16.$ (37)
The masses obtained here, using Eq. (36), are presented in Table 4, with respectives uncertainties, using the mass of the state $\rho(770)$ as an input from PDG. The Regge trajectory for these mesons is given by
$J=(0.87\pm 0.01)M^{2}+0.51\pm 0.02,$ (38)
with $\chi^{2}$/ndf$=8.68/4=2.17$ and is presented in the upper left panel of Fig. 2 as the blue line. The red dots represent these masses with their uncertainties.
In order to show that the linear AHW model indeed present asymptotic linear Regge trajectories we also plot $J\times M^{2}$ for these mesons with even $\mathcal{CP}$, isospin $I=1$, and $S\geq 1$, for the first 50 states. The trajectory in this case becomes
$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01,$ (39)
with $\chi^{2}$/ndf$=72.96/48=1.52.$ This trajectory is presented in the upper left panel of Fig. 3. Both trajectories can be compared with the one obtained from PDG data shown in Table 2 and are in good agreement with it.
Table 4: Masses in MeV for mesons classified under even or odd $\mathcal{CP}$ with isospin $I=0,1$ and spins $S\geq 1$ from the Linear AHW model, determined from Eq. (36), and the percentual deviations $\delta_{\rm Lin}$ of the masses from this model compared with PDG data. For each family the mass of the state with $J=S=1$ is taken as an input. In the last column we show the total dimension $\Delta_{\rm Lin}$ for each state determined by Eq. (35).
$J^{PC}$ $\mathcal{CP}$ I Meson Linear AHW $\delta_{\rm LAHW}$ $\Delta_{\rm Lin}$
$1^{--}$ even 1 $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
$2^{++}$ even 1 $a_{2}(1320)$ $1296.75\pm 21$ $1.53\%$ $1.85$
$3^{--}$ even 1 $\rho_{3}(1690)$ $1678.61\pm 30$ $0.60\%$ $3.03$
$4^{++}$ even 1 $a_{4}(1970)$ $1993.84\pm 38$ $1.36\%$ $3.87$
$5^{--}$ even 1 $\rho_{5}(2350)$ $2268.03\pm 44$ $2.66\%$ $4.49$
$6^{++}$ even 1 $a_{6}(2450)$ $2513.71\pm 51$ $2.60\%$ $4.96$
$1^{--}$ even 0 $\omega(782)$ $782.66\pm 7.83$ $0\%$ 00
$2^{++}$ even 0 $f_{2}(1270)$ $1294.28\pm 16$ $1.47\%$ $1.76$
$3^{--}$ even 0 $\omega_{3}(1670)$ $1669.09\pm 22$ $0.13\%$ $2.88$
$4^{++}$ even 0 $f_{4}(2050)$ $1978.53\pm 28$ $1.96\%$ $3.67$
$5^{--}$ even 0 $\omega_{5}(2250)$ $2247.67\pm 33$ $0.10\%$ $4.24$
$6^{++}$ even 0 $f_{6}(2510)$ $2488.81\pm 38$ $0.97\%$ $4.67$
$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 5.0$ $0\%$ 00
$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1667.30\pm 7.89$ $0.20\%$ $0.48$
$3^{+-}$ odd 1 $b_{3}(2030)$ $2000.60\pm 10.75$ $1.55\%$ $0.65$
$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2281.03\pm 13.31$ $1.38\%$ $0.65$
$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 10$ $0\%$ 00
$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1646.87\pm 17$ $1.85\%$ $0.72$
$3^{+-}$ odd 0 $h_{3}(2025)$ $2009.96\pm 24$ $0.74\%$ $1.07$
$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2314.19\pm 31$ $0.59\%$ $1.22$
Figure 2: Regge trajectories (blue lines) for mesons, with masses $m$ expressed in GeV, from the Linear AHW model, with total dimension given by Eq. (35). Red dots are the values obtained from Eq. (36) with uncertainties and the blue lines are obtained by linear regression of these points, for each meson family. Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
Figure 3: Regge trajectories for mesons, with masses $m$ expressed in GeV, from the linear AHW model, with total dimension given by Eq. (35). Red dots are the values obtained from Eq. (36) with uncertainties and the blue lines are the Regge trajectories, obtained by linear regression of these 50 points, for each meson family. Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
IV.2 Linear AHW for mesons with even ${\cal CP}$, $I=0$, and $S\geq 1$
For mesons in the linear AHW model with even ${\cal CP}$, $I=1$, and $S\geq 1$, the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=6.67\pm 0.13$ and $\epsilon=0$. So, the total dimension reads
$\Delta_{\rm Lin}=(6.67\pm 0.13)\sqrt{S}-2.67\pm 0.13.$ (40)
The masses obtained by this model, using Eq. (36), are presented with respectives uncertainties in Table 4, taking the mass of the state $\omega(782)$ as an input from PDG. Furthermore, we also present in this table the total dimension and the percentual deviations from PDG data for each mesonic state. The Regge trajectory obtained by linear regression of these masses (first six states) is presented in the upper right panel of Fig. 2 and is given by
$J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02,$ (41)
with $\chi^{2}$/ndf$=10.12/4=2.53.$ The upper right panel of Fig. 3 presents the Regge trajectory for the first 50 sates with even ${\cal CP}$, $I=0$, and $S\geq 1$, and it is given by
$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01,$ (42)
with $\chi^{2}$/ndf$=32.16/48=0.67.$ Both equations can be compared with the one obtained from PDG data, shown in Table 2, and are in good agreement with it.
IV.3 Linear AHW for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$
For mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$ ($b_{1}(1235)$ family), the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=3.19\pm 0.03$ and $\epsilon=0.05$, so the total dimension becomes
$\Delta_{\rm Lin}=(3.19\pm 0.03)S^{0.55}+0.81\pm 0.03.$ (43)
The masses obtained in this model, using Eq. (36), are presented in Table 4. The Regge trajectory obtained for the first four points is presented in Fig. 2 and is given by
$J=(0.81\pm 0.01)M^{2}-0.24\pm 0.03,$ (44)
with $\chi^{2}$/ndf$=2.86/2=1.43.$
For 50 points the trajectory is in Fig. 3 and is given by
$J=(0.8549\pm 0.0003)M^{2}-0.42\pm 0.01,$ (45)
with $\chi^{2}$/ndf$=163/48=3.39.$ Both equations, considering the uncertainties, are in agreement with the Regge trajectory obtained from PDG data, shown in Table 2.
IV.4 Linear AHW for mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$
The parameters in Eq. (35) that minimizes $Q$ for meson with mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$ ($h_{1}(1170)$ family) are $a_{\rm Lin}=3.79\pm 0.08$ and $\epsilon=0.04$, so the total dimension is
$\Delta_{\rm Lin}=(3.79\pm 0.08)S^{0.54}+0.21\pm 0.08.$ (46)
The masses obtained in this model, using Eq. (36), are presented in Table 4. The Regge trajectory for the first four points is presented in Fig. 2 and is
$J=(0.7509\pm 0.033)M^{2}-0.03\pm 0.01,$ (47)
with $\chi^{2}$/ndf$=0.22/2=0.11.$ The Fig. 3 presents the Regge trajectory for the first 50 points, given by
$J=(0.7667\pm 0.0003)M^{2}-0.06\pm 0.01,$ (48)
with $\chi^{2}$/ndf$=2.20/48=0.046.$ Both trajectories are in good agreement the one from PDG shown in Table 2.
V Conclusions
In this work, we proposed anomalous and linear HW models for light unflavored mesons. We have shown that the masses and Regge trajectories produced by these models are better than that obtained from the original HW and SW models.
In particular, the linear AHW model produce asymptotically linear Regge trajectories, circumventing the well-kown problem of non-linearity of the original HW model.
Finally, we expect to succesfully extend the analysis presented here to other particles as heavy and heavy-light mesons, as well as baryons. This is presently under investigation.
Acknowledgments
R.A.C.-S. is supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) under finance code 001. H.B.-F. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under grant ##\## 310346/2023-1.

### v2

@@
 Abstract
-In this work we consider anomalous and linear holographic hard wall (HW) models for light unflavored mesons inspired by the AdS/CFT correspondence.
-The anomalous dimensions depend on the logarithm of the spin $S$ of the meson state and come from a semiclassical analysis of gauge/string duality.
-The anomalous HW model produces very good masses and good Regge trajectories for mesons compared with PDG data.
-Inspired by this anomalous HW model we also propose a phenomenological modification of the dimension of the boundary operators such that the model produces asymptotic linear Regge trajectories.
+[…]
 I Introduction
 The AdS/CFT correspondence Aharony:1999ti establishes relations between Yang-Mills (YM) and higher dimensional gravitational theories.
-One of the amazing properties of the correspondence is that if YM theories are strongly coupled, the gravitational one is weak coupled allowing for the solution of a myriad of phenomena from the quark-gluon plasma to condensed matter (for reviews, see e.g., Casalderrey-Solana:2011dxg ; Zaanen:2015oix ).
-In particular, hadronic spectra were described by different proposals as the holographic hard wall (HW) and the soft wall (SW) models Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us .
+One of the amazing properties of this correspondence is that if the YM theories are strongly coupled, the dual gravitational one is weak coupled allowing for the solution of a myriad of phenomena from the quark-gluon plasma to condensed matter (for reviews, see e.g., Casalderrey-Solana:2011dxg ; Zaanen:2015oix ).
+In particular, hadronic spectra were described by different proposals such as the holographic hard wall (HW) and the soft wall (SW) models Polchinski:2001tt ; Polchinski:2002jw ; Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us .
 The HW model Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct is well known for producing reasonable hadronic masses but non-linear (parabolic) Regge trajectories.
 On the other side, the SW model Karch:2006pv ; Colangelo:2008us also give good masses with the advantage of generating linear Regge trajectories.
-In this work, we are going to consider the inclusion of anomalous dimensions for the boundary operators describing light unflavored mesons with spins $S\geq 1$.
-These anomalous dimensions depend on the logarithm of the spin $S$ of the mesonic state inspired by a semiclasical analysis of gauge/string dualities Gubser:2002tv and was used recently to discuss the case of even spins glueballs Costa-Silva:2023vuu .
+For other interesting QCD properties described by the HW model, see e.g.
+Boschi-Filho:2005nmp ; Boschi-Filho:2006hfm ; Herzog:2006ra ; BallonBayona:2007vp ; Grigoryan:2007my ; Kwee:2007dd ; Grigoryan:2007wn ; Jo:2009xr ; Costa:2013uia ; Craps:2013iaa ; Mamo:2015dea ; Ballon-Bayona:2017bwk .
+In this work, we are going to consider the inclusion of anomalous dimensions for the boundary operators describing light unflavored mesons with spins $S\geq 1$ in the HW model.
+These anomalous dimensions depend on the logarithm of the spin $S$ of the mesonic state inspired by a semiclasical analysis of gauge/string dualities Gubser:2002tv and was used recently to discuss the case of even spin glueballs Costa-Silva:2023vuu .
 We find very good masses and Regge trajectories for the studied mesons when compared with PDG data PDG:2022pth , better than the ones from the HW and SW models.
-Furthermore, we discuss a phenomenological modification of the dimension of the boundary operators representing the mesonic states in order to obtain asymptotic linear Regge trajectories, similar to the case of glueballs.
+Furthermore, we discuss a phenomenological modification of the dimension of the boundary operators representing the mesonic states in order to obtain asymptotically linear Regge trajectories, similar to the case of glueballs.
 Actually, the linearization of the mesonic states require an additional parameter to obtain the best fit to PDG data.
@@
 Chen:2021kfw .
+For later reference, we present in Table 1 the masses of the meson families starting with the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$, with their corresponding Regge trajectories.
+These families are classified according to their $PC$ parity (even when $P$ and $C$ have the same sign or odd with opposite signs) and isospins $I=0,1$.
+Note that, according to PDG notation, mesons are classified as $J^{PC}$, where $J$ is the total angular momenta, $P$ is the parity and $C$ the charge conjugation of the state.
+In our notation, we simply call the total angular momenta as the “spin” $S$, so that $J=S$.
+The $G$ parity ($G=(-)^{I+S}$) of these mesons is also included in this table following PDG notation but is not relevant in the present analysis.
+$I^{G}(J^{PC})$ Meson PDG (MeV) PDG Regge trajetories (GeV)
+$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 0.25$
+$1^{-}(2^{++})$ $a_{2}(1320)$ $1316.9\pm 0.9$
+$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1688.8\pm 2.1$ $J=(0.88\pm 0.04)M^{2}+0.47\pm 0.15$
+$1^{-}(4^{++})$ $a_{4}(1970)$ $1967\pm 16$
+$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2330\pm 35$
+$1^{-}(6^{++})$ $a_{6}(2450)$ $2450\pm 130$
+$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 0.13$
+$0^{+}(2^{++})$ $f_{2}(1270)$ $1275.5\pm 0.8$
+$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1667\pm 4$ $J=(0.90\pm 0.02)M^{2}+0.47\pm 0.07$
+$0^{+}(4^{++})$ $f_{4}(2050)$ $2018\pm 11$
+$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2250\pm 70$
+$0^{+}(6^{++})$ $f_{6}(2510)$ $2465\pm 50$
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 3.2$
+$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1670.6\pm 2.05$ $J=(0.83\pm 0.04)M^{2}-0.30\pm 0.16$
+$1^{+}(3^{+-})$ $b_{3}(2030)$ $2032\pm 12$
+$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2250\pm 15$
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 6$
+$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1617\pm 5$ $J=(0.73\pm 0.02)M^{2}+0.03\pm 0.06$
+$0^{-}(3^{+-})$ $h_{3}(2025)$ $2025\pm 20$
+$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2328\pm 38$
+Table 1: Masses in MeV for light unflavored mesons classified under even or odd $PC$, with isospins $I=0,1$, and spins $J=S\geq 1$ from PDG PDG:2022pth , and their corresponding Regge trajectories in GeV.
 This work is organized as follows.
 In Section II, we review the HW model for light unflavored mesons and in Section III we discuss the introduction of anomalous dimensions in the HW model and obtain the spectra for the mesonic families starting with the states of the $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$.
-Then, in Section IV we modify the dimension of the mesonic boundary operators such that we find asymptotic linear Regge trajectories for the mesons already discussed in the previous section.
+Then, in Section IV we modify the dimension of the mesonic boundary operators such that we find asymptotically linear Regge trajectories for the mesons already discussed in the previous section.
 Finally, in Section V we present our conclusions.
@@
 In this work we are interested in studying the spectra of light unflavored mesons, and their Regge trajectories using anomalous and linear holographic HW models.
-These mesons can be classified in four families according to their behavior under $\mathcal{CP}$ transformations (even or odd), isospin $I=0$ and $1$, and spins $S\geq 1$.
+These mesons can be classified in four families according to their behavior under ${PC}$ transformations (even or odd), isospin $I=0$ or $1$, and spins $S\geq 1$.
 These four families imply four distinct Regge trajectories.
-In order to describe such objects using the holographic HW model, we need to study a massive vector field in $AdS_{5}$. We start with the action
+In order to describe such objects using the holographic HW model, we need to study a massive vector field in $AdS_{5}$. We start with the action Aharony:1999ti
 $S=-\frac{1}{2}\int d^{5}x\sqrt{-g}\bigg{[}\frac{1}{2}g^{pm}g^{qn}F_{mn}F_{pq}+M_{5}^{2}g^{pm}A_{p}A_{m}\bigg{]},$ (1)
-where $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$ is the stress tensor of the Proca field with mass $M_{5}$ in the five-dimensional AdS space.
+where $F_{mn}=\partial_{m}A_{n}-\partial_{n}A_{m}$ is the stress tensor of the Proca field $A_{m}$ with mass $M_{5}$ in the five-dimensional AdS space.
 Moreover, $g$ is the determinant of the metric of $AdS_{5}$,
@@
 $z^{2}\frac{d^{2}\psi(z)}{dz^{2}}+z\frac{d\psi(z)}{dz}-q^{2}z^{2}\psi(z)-[1+(M_{5}R)^{2}]f(z)=0.$ (6)
-This is Bessel equation, whose solutions are linear combination of Bessel ($J_{\nu}$) and Neumann ($Y_{\nu})$ functions of order $\nu=\sqrt{1+(M_{5}R)^{2}}$.
+This is Bessel equation, whose solutions are linear combinations of Bessel ($J_{\nu}$) and Neumann ($Y_{\nu})$ functions both of order $\nu=\sqrt{1+(M_{5}R)^{2}}$.
 Since we are interested only in regular solutions in the bulk, we will consider just the Bessel solution
@@
 (13) below).
-On the other hand, the mesons with spin $S$ with $S=1,2,3,...$, that we are interested in are described by operators in the boundary Minkowski space as Gubser:2002tv  [^1] It is also possible to alternatively write the mesons with spin $S$ as deTeramond:2005su ${\cal O}^{\mu\,\mu_{1}...\mu_{S}}_{2+S}=\;\bar{\psi}\gamma^{\mu}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (8) such that $\Delta=2+S.$ (9) However, in this case, the results for the AHW and LAHW are not so good as the ones discussed in Sections III and IV, with $\Delta=3+S$.
-${\cal O}^{\mu_{1}...\mu_{S}}_{3+S}=\;\bar{\psi}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (10)
-where $\psi$ and $\bar{\psi}$ represent quark and antiquark fields while $D_{\mu_{i}}$ with $i=1,\dots,S$ are symmetrized covariant derivatives.
+On the other hand, the mesons with spin $S=1,2,3,...$, that we are interested in are described by operators in the boundary Minkowski space as Gubser:2002tv
+${\cal O}_{\mu_{1}...\mu_{S}}^{3+S}=\;\bar{\psi}D_{\{\mu_{1}\dots}D_{\mu_{S}\}}\psi\;;\qquad(S=1,2,3,...)\,,$ (8)
+where $\psi$ and $\bar{\psi}$ represent quark and antiquark fields while $D_{\mu_{i}}$ with $i=1,\dots,S$, are symmetrized covariant derivatives Operador .
 Then, the canonical dimensions of these operators are given by
 $\Delta=\Delta_{\rm can.}=3+S.$ (9)
-The link between the operators $\cal O$ and string excitations at $AdS_{5}$ can be constructed from its conformal dimension.
-From the AdS/CFT correspondence, the vector field in $AdS_{5}$ that we are describing is a $1$-form, so its holographic mass $M_{5}$ is related to the conformal dimension $\Delta$ of boundary operators $\cal O$, is given by
+The link between the operators $\cal O$ and string excitations in $AdS_{5}$ can be constructed from its conformal dimension.
+From the AdS/CFT correspondence, the vector field in $AdS_{5}$ that we are describing is a $1$-form, so its holographic mass $M_{5}$ is related to the conformal dimension $\Delta$ of boundary operators $\cal O$, and given by Aharony:1999ti
 $(M_{5}R)^{2}=(\Delta-1)(\Delta-3).$ (10)
@@
 $\nu=S+1.$ (12)
-The HW model consists in the holographic relation between fields in AdS space with an IR cutoff that breaks the conformal symmetry of the boundary theory.
+The HW model consists in the holographic relation between bulk fields in AdS space with an infrared (IR) cutoff that breaks the conformal symmetry of the boundary theory.
 That means a hard cutoff in the holographic $z$ coordinate, so we are working in an AdS slice, given by $0\leq z\leq z_{\rm max}$, where
-$z_{\rm max}=\frac{1}{\Lambda_{\rm QCD}},$ (15)
-with $\Lambda_{\rm QCD}$ a convenient energy scale related to be fixed later. Then the solution of the vector bulk field, Eq. (7), assumes the form
+$z_{\rm max}=\frac{1}{\Lambda_{\rm HW}},$ (13)
+with $\Lambda_{\rm HW}$ a convenient energy scale in the HW model to be fixed later (see Eq. (17) below). Then, the solution of the vector bulk field, Eq. (7), assumes the form
 $f(z)\sim J_{\nu}(m_{\nu,k}\,z),$ (14)
 where $m_{\nu,k}$ is identified with the mass of the boundary operator $\cal O$ that lives in the four-dimensional Minkowski space, describing mesons.
-Note that $k=1,2,3...$ denote radial excitations, where $k=1$ is the corresponding ground state.
+Note that $k=1,2,3...,$ denote radial excitations, where $k=1$ is the corresponding ground state.
 In this work we are interested in angular momenta excitations, so we take $k=1$ for all the objects from now on.
+$I^{G}(J^{PC})$ Meson HW $\delta_{\rm HW}$ $\Lambda_{\rm HW}$ HW Regge Trajectories (GeV) SW $\delta_{\rm SW}$
+$1^{+}(1^{--})$ $\rho(770)$ $775.26$ $0\%$ $775.26$ $0\%$
+$1^{-}(2^{++})$ $a_{2}(1320)$ $963.13$ $25.9\%$ $1096.38$ $16.7\%$
+$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1145.52$ $32.2\%$ 150.96 $J=(2.25\pm 0.10)M^{2}-0.12\pm 0.19$ $1342.79$ $20.5\%$
+$1^{-}(4^{++})$ $a_{4}(1970)$ $1324.12$ $32.7\%$ $1550.52$ $21.2\%$
+$1^{+}(5^{--})$ $\rho_{5}(2350)$ $1499.93$ $35.6\%$ $1733.53$ $25.6\%$
+$1^{-}(6^{++})$ $a_{6}(2450)$ $1673.57$ $31.7\%$ $1898.99$ $22.5\%$
+$0^{-}(1^{--})$ $\omega(782)$ $782.66$ $0\%$ $782.66$ $0\%$
+$0^{+}(2^{++})$ $f_{2}(1270)$ $972.31$ $23.8\%$ $1106.85$ $13.2\%$
+$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1156.44$ $30.6\%$ 152.24 $J=(2.21\pm 0.10)M^{2}-0.12\pm 0.19$ $1355.61$ $18.7\%$
+$0^{+}(4^{++})$ $f_{4}(2050)$ $1336.74$ $33.8\%$ $1565.32$ $22.4\%$
+$0^{-}(5^{--})$ $\omega_{5}(2250)$ $1514.23$ $32.7\%$ $1750.08$ $22.2\%$
+$0^{+}(6^{++})$ $f_{6}(2510)$ $1689.52$ $31.5\%$ $1917.12$ $22.2\%$
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5$ $0\%$ $1229.5$ $0\%$
+$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1527.41$ $8.57\%$ 239.41 $J=(1.03\pm 0.05)M^{2}-0.48\pm 0.15$ $1738.78$ $4.08\%$
+$1^{+}(3^{+-})$ $b_{3}(2030)$ $1816.70$ $10.6\%$ $2129.56$ $4.80\%$
+$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2099.95$ $6.67\%$ $2459.00$ $9.29\%$
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166.0$ $0\%$ $1166$ $0\%$
+$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1448.56$ $10.4\%$ 227.04 $J=(1.15\pm 0.05)M^{2}-0.48\pm 0.15$ $1648.97$ $1.98\%$
+$0^{-}(3^{+-})$ $h_{3}(2025)$ $1722.87$ $14.9\%$ $2019.57$ $0.27\%$
+$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $1991.49$ $14.5\%$ $2332.00$ $0.17\%$
+Table 2: Masses in MeV for light unflavored mesons classified under even or odd $PC$, with isospins $I=0,1$, and spins $J=S\geq 1$ from the HW model, Eq.
+(16), and their corresponding Regge trajectories in GeV, and $\Lambda_{\rm HW}$ in MeV for each family.
+For comparison, we also show masses from the SW model, according to $m_{S}^{2}\propto S$ Karch:2006pv .
+The percentage deviations of both models refer to PDG data presented in Table 1.
 Imposing Dirichlet boundary conditions (for $k=1$) in Eq. (14) at $z=z_{\rm max}$, one obtains
-$J_{\nu}\bigg{(}\frac{m_{\nu,1}}{\Lambda_{\rm QCD}}\bigg{)}=0.$ (17)
+$J_{\nu}\bigg{(}\frac{m_{\nu,1}}{\Lambda_{\rm HW}}\bigg{)}=0.$ (15)
 Defining the first zero $\chi_{\nu,1}$ of the Bessel function of order $\nu$, such that $J_{\nu}(\chi_{\nu,1})=0$, and making use of the relation Eq.
 (12), we can write the mass of the mesonic state with spin $S$ as
-$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=S+1).$ (18)
+$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm HW};\qquad\qquad(\nu=S+1).$ (16)
 In particular, from the state with $S=1$ we fix
-$\Lambda_{\rm QCD}=\frac{m_{2,1}}{\chi_{2,1}},$ (19)
-taking $m_{2,1}$ as an input from PDG for the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$.
-The corresponding values of $\Lambda_{\rm QCD}$ found here for each mesonic family are shown in Table 2.
+$\Lambda_{\rm HW}=\frac{m_{2,1}}{\chi_{2,1}},$ (17)
+taking $m_{2,1}$ as an input from PDG PDG:2022pth for the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$.
 The meson masses obtained from the HW model, using Eq.
-(18), are shown in Table 1, together with the SW model predictions with the relative deviations with respect to PDG data PDG:2022pth .
-In Table 2, we also show the equations for the Regge trajectories found from PDG data for the same mesonic families.
-The process described above is known as the holographic HW model and was used in the past to obtain hadronic spectra among other properties.
-It is a well-known fact that this model implies non-linear Regge trajectories.
-In the next section we develop an improvement of the HW model, by considering anomalous dimensions, in order to obtain better masses compared with the PDG ones and better Regge trajectories for these meson families.
-Table 1: Masses in MeV for light unflavored mesons classified under even or odd $\mathcal{CP}$ and with isospin $I=0,1$, and spins $S\geq 1$ from PDG PDG:2022pth , from the HW Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct and SW Karch:2006pv ; Colangelo:2008us models, and the percentual deviations $\delta_{\rm HW}$ and $\delta_{\rm SW}$ of the masses from these two models compared with PDG data.
-The HW masses are determined from Eq.
-(18), using the values of $\Lambda_{\rm QCD}$ for each family (same $\mathcal{CP}$ parity and Isospin) shown in Table 2, and the masses from the SW model come from the relation $m_{S}^{2}\sim S$ Karch:2006pv .
-The mass of the state with $J=S=1$ for each family is used as an input for the HW and SW models.
-$J^{PC}$ ${CP}$ $I$ Meson PDG HW $\delta_{\rm HW}$ SW $\delta_{\rm SW}$
-$1^{--}$ even 1 $\rho(770)$ $775.26\pm 0.25$ $775.26$ $0\%$ $775.26$ $0\%$
-$2^{++}$ even 1 $a_{2}(1320)$ $1316.9\pm 0.9$ $963.13$ $25.9\%$ $1096.38$ $16.7\%$
-$3^{--}$ even 1 $\rho_{3}(1690)$ $1688.8\pm 2.1$ $1145.52$ $32.2\%$ $1342.79$ $20.5\%$
-$4^{++}$ even 1 $a_{4}(1970)$ $1967\pm 16$ $1324.12$ $32.7\%$ $1550.52$ $21.2\%$
-$5^{--}$ even 1 $\rho_{5}(2350)$ $2330\pm 35$ $1499.93$ $35.6\%$ $1733.53$ $25.6\%$
-$6^{++}$ even 1 $a_{6}(2450)$ $2450\pm 130$ $1673.57$ $31.7\%$ $1898.99$ $22.5\%$
-$1^{--}$ even 0 $\omega(782)$ $782.66\pm 0.13$ $782.66$ $0\%$ $782.66$ $0\%$
-$2^{++}$ even 0 $f_{2}(1270)$ $1275.5\pm 0.8$ $972.31$ $23.8\%$ $1106.85$ $13.2\%$
-$3^{--}$ even 0 $\omega_{3}(1670)$ $1667\pm 4$ $1156.44$ $30.6\%$ $1355.61$ $18.7\%$
-$4^{++}$ even 0 $f_{4}(2050)$ $2018\pm 11$ $1336.74$ $33.8\%$ $1565.32$ $22.4\%$
-$5^{--}$ even 0 $\omega_{5}(2250)$ $2250\pm 70$ $1514.23$ $32.7\%$ $1750.08$ $22.2\%$
-$6^{++}$ even 0 $f_{6}(2510)$ $2465\pm 50$ $1689.52$ $31.5\%$ $1917.12$ $22.2\%$
-$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 3.2$ $1229.5$ $0\%$ $1229.5$ $0\%$
-$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1670.6\pm 2.05$ $1527.41$ $8.57\%$ $1738.78$ $4.08\%$
-$3^{+-}$ odd 1 $b_{3}(2030)$ $2032\pm 12$ $1816.70$ $10.6\%$ $2129.56$ $4.80\%$
-$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2250\pm 15$ $2099.95$ $6.67\%$ $2459.00$ $9.29\%$
-$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 6$ $1166$ $0\%$ $1166$ $0\%$
-$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1617\pm 5$ $1352$ $16.4\%$ $1648.97$ $1.98\%$
-$3^{+-}$ odd 0 $h_{3}(2025)$ $2025\pm 20$ $1709$ $15.6\%$ $2019.57$ $0.27\%$
-$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2328\pm 38$ $1882$ $19.2\%$ $2332.00$ $0.17\%$
-Table 2: The Regge trajectory in GeV units for each light unflavored meson family (same $\mathcal{CP}$ parity and same isospin $I$) from the best fit of PDG data shown in Table 1.
-We also show the mass of the state with $J=S=1$ for each family, and the corresponding $\Lambda_{\rm QCD}$ value, both in MeV, determined from Eq.
-(19).
-$J^{PC}$ ${CP}$ $I$ Meson PDG (MeV) PDG Regge Trajectory (GeV) $\Lambda_{\rm QCD}$ (MeV)
-$1^{--}$ even 1 $\rho(770)$ $775.26\pm 0.25$ $J=(0.88\pm 0.04)M^{2}+0.47\pm 0.15.$ $150.96$
-$1^{--}$ even 0 $\omega(782)$ $782.66\pm 0.13$ $J=(0.90\pm 0.02)M^{2}+0.47\pm 0.07$ 152.24
-$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 3.2$ $J=(0.83\pm 0.04)M^{2}-0.30\pm 0.16$ 239.41
-$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 6$ $J=(0.73\pm 0.02)M^{2}+0.03\pm 0.06$ 227.04
+(16), are shown in Table 2, with the corresponding values of $\Lambda_{\rm HW}$ and Regge trajectories found for each mesonic family, together with the SW model predictions with the relative deviations with respect to PDG data, presented in Table 1.
+Figure 1: Regge trajectories ($J\times M^{2}$) from the HW model (blue lines) and from PDG data (orange lines) for mesons, with masses expressed in GeV.
+Red dots are the values obtained from Eq.
+(16) in the HW model with uncertainties shown in Table 2, while the black dots correspond to PDG data shown in Table 1.
+The blue and orange lines are obtained by linear regression of these points.
+Upper left panel: mesons with even ${PC}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
+The process described above is known as the holographic HW model and was used in the past to obtain hadronic spectra among other properties Polchinski:2001tt ; Polchinski:2002jw ; Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Boschi-Filho:2005nmp ; Boschi-Filho:2006hfm ; Herzog:2006ra ; BallonBayona:2007vp ; Grigoryan:2007my ; Kwee:2007dd ; Grigoryan:2007wn ; Jo:2009xr ; Costa:2013uia ; Craps:2013iaa ; Mamo:2015dea ; Ballon-Bayona:2017bwk .
+Nevertheless, it is a well-known fact that this model implies non-linear Regge trajectories Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv in contrast with experimental data (see, e.g., Spector:1967xsn ; Kanki:1975fp )
+For comparison, we plot in Figure 1 experimental data (from Table 1) for the meson families starting with the states $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$ together with HW masses (shown in Table 2) with uncertainties and the corresponding Regge trajectories obtained by linear regression of these points.
+The equations for these trajectories are also shown in Tables 1 and 2.
+As one can see from this Figure, the agreement between HW and PDG masses is not very good, specially for the $\rho(770)$ and $\omega(782)$ families.
+In the next section we develop an improvement of the HW model, by considering anomalous dimensions, in order to obtain better masses compared with the PDG ones, and Regge trajectories for these meson families which are closer to experimental data.
 III Anomalous HW model
-In an interacting theory, the dimension of an operator is no longer canonical, thanks to renormalization that implies in anomalous contributions ($\Delta_{\rm Anom}$).
+In an interacting theory, the dimension of an operator is no longer canonical, due to renormalization that introduces anomalous contributions ($\Delta_{\rm Anom}$).
 So, the total dimension of an operator is given by its canonical plus the anomalous dimensions:
 $\Delta_{\rm Total}=\Delta_{\rm can.}+\Delta_{\rm Anom}\,.$ (18)
-From a well known semi-classical gauge/string duality analysis Gubser:2002tv , it is found that the anomalous dimension increases logarithmically with the spín $S$ of the boundary operators as
+From a well known semi-classical gauge/string duality analysis Gubser:2002tv , it is found that the anomalous dimension increases logarithmically with the spin $S$ of the boundary operators as
 $\Delta_{\rm Anom}=\frac{\sqrt{\lambda}}{\pi}\ln{\bigg{(}\frac{S}{\sqrt{\lambda}}\bigg{)}}+\mathcal{O}(S^{0}),$ (19)
-where $\lambda$ is ’t Hooft coupling. Inspired by this result, we consider an anomalous dimension given by
-$\Delta_{\rm Log}=a\ln(S),$ (22)
-where $a$ is a constant that will be determined by best fit for each meson family.
+where $\lambda$ is the ’t Hooft coupling. Inspired by this result, we consider an anomalous dimension given by
+$\Delta_{\rm Log}=a_{\rm Log}\ln(S),$ (20)
+where $a_{\rm Log}$ is a constant that will be determined by best fit for each meson family.
 Then, the anomalous HW (AHW) model consists in considering bulk fields in AdS slice, Eq.
@@
 (18) such that the mass associated with boundary operators, now takes the form
-$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=S+1+\Delta_{\rm Log}),$ (23)
+$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm HW};\qquad\qquad(\nu=S+1+\Delta_{\rm Log}),$ (21)
 where we used the fact that $\Delta_{\rm can.}=3+S$ for these mesonic states and that the order of the corresponding Bessel function is $\nu=\Delta_{\rm Total}-2$.
-The value of $\Lambda_{\rm QCD}$ is determined for each meson familiy using as inputs the masses of the states with $S=1$ from PDG, namely $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$.
-These values of $\Lambda_{\rm QCD}$ coincide with the ones used in the usual HW model, listed in Table 2, once Eq.
-(22) implies a vanishing anomalous dimension for $S=1$.
-In order to obtain the best numerical fit, let us define the percentual deviation:
-$q_{i}=\frac{|m_{i}-M_{i}|}{M_{i}},$ (24)
+The value of $\Lambda_{\rm HW}$ is determined for each meson familiy using as inputs the masses of the states with $S=1$ from PDG, namely $\rho(770)$, $\omega(782)$, $b_{1}(1235)$, and $h_{1}(1170)$.
+By construction, these values of $\Lambda_{\rm HW}$ are the same as those ones used in the usual HW model, listed in Table 2, since Eq.
+(20) implies a vanishing anomalous dimension $\Delta_{\rm Log}$ for $S=1$.
+In order to obtain the best numerical fit, let us define the fractional deviation:
+$\delta_{i}=\frac{|m_{i}-M_{i}|}{M_{i}},$ (22)
 where $m_{i}$ is the mass of a determined particle obtained by our model, from Eq. (21), and $M_{i}$ its mass obtained from PDG, listed in Table 1. Further, we define the function $Q$ as
-$Q=\sqrt{\sum_{i}q_{i}^{2}}.$ (25)
-So, the best fit method adopted here corresponds to minimize the function $Q$ for each meson family in the AHW model.
-We will also consider an uncertainty for the parameter $a$, defined in Eq. (22), which implies in uncertainties for the masses ${m_{i}}$, obtained by error propagation, accordingly to the formula:
+$Q=\sqrt{\sum_{i}\delta_{i}^{2}}.$ (23)
+So, the best fit method adopted here corresponds to minimization of the deviations from PDG data given by the function $Q$ for each meson family in the AHW model.
+$I^{G}(J^{PC})$ Meson AHW Log $\delta_{\rm Log}$ $\Delta_{\rm Log}$
+$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
+$1^{-}(2^{++})$ $a_{2}(1320)$ $1309.95\pm 22$ $0.53\%$ $1.92$
+$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1681.02\pm 31$ $0.46\%$ $3.04$
+$1^{-}(4^{++})$ $a_{4}(1970)$ $1988.75\pm 37$ $1.11\%$ $3.84$
+$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2262.37\pm 43$ $2.90\%$ $4.46$
+$1^{-}(6^{++})$ $a_{6}(2450)$ $2514.58\pm 47$ $2.64\%$ $4.96$
+$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
+$0^{+}(2^{++})$ $f_{2}(1270)$ $1302.58\pm 22$ $2.12\%$ $1.81$
+$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1666.4\pm 30$ $0.04\%$ $2.87$
+$0^{+}(4^{++})$ $f_{4}(2050)$ $1969.63\pm 37$ $2.40\%$ $3.62$
+$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2240.19\pm 42$ $0.44\%$ $4.20$
+$0^{+}(6^{++})$ $f_{6}(2510)$ $2490.23\pm 46$ $1.02\%$ $4.68$
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
+$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1640.57\pm 27$ $1.80\%$ $0.39$
+$1^{+}(3^{+-})$ $b_{3}(2030)$ $1991.55\pm 37$ $1.99\%$ $0.62$
+$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2316.73\pm 44$ $2.97\%$ $0.78$
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
+$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1629.93\pm 25$ $0.80\%$ $0.66$
+$0^{-}(3^{+-})$ $h_{3}(2025)$ $2003.12\pm 33$ $1.08\%$ $1.04$
+$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2339.01\pm 39$ $0.47\%$ $1.32$
+Table 3: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the AHW Log model, determined from Eq.
+(21), and the percentage deviations $\delta_{\rm Log}=\delta_{i}\times 100\%$, with $\delta_{i}$ given by Eq.
+(22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $Q$, Eq.
+(23).
+For each family the mass of the state with $J=S=1$ is taken as an input from PDG.
+In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq.
+(20).
+Further, we want to know by how much these masses are close to the linear Regge trajectory, obtained from the model, for each family. In order to measure this, we compute
+$R^{2}=1-\frac{\sum_{i}[m_{i}-\mathcal{M}_{i}(S)]^{2}}{\sum_{i}(m_{i}-\bar{m})^{2}},$ (24)
+where $m_{i}$ is the mass obtained from the AHW model, Eq.
+(21), for each state $i$, while ${\cal M}_{i}(S)$ is the corresponding mass of this state with angular momentum $S$, computed from the equation of the Regge trajectory and $\bar{m}$ is the average of the masses $m_{i}$.
+In a perfect model one would have $R^{2}=1$, so the closer $R^{2}$ is to one, the best is the match between the output masses of the model and the linear Regge trajectory.
+We also consider an uncertainty for the parameter $a_{\rm Log}$, defined in Eq. (20), which implies uncertainties for the masses ${m_{i}}$, obtained by error propagation, according to the formula:
 $\delta{m_{i}}=\sqrt{\left(\frac{\chi_{\nu_{i},1}}{\chi_{2,1}}\right)^{2}\left(\delta m_{0}\right)^{2}+\left(\frac{m_{0}}{\chi_{2,1}}\right)^{2}\left(\delta\chi_{\nu_{i},1}\right)^{2}},$ (25)
 where $\delta m_{0}$ and $\delta\chi_{\nu_{i},1}$ are the uncertainties in the mass of the state with $S=1$ and in the zero of Bessel function that is related to the mass $m_{i}$.
-The uncertainty $\delta m_{0}$ of the state with $S=1$ for each meson family in our model will be the arithmetic average of the percentual uncertainties of PDG masses $M_{i}$ for that given family.
-Defined our criterion for best fit and how we compute the uncertainties $\delta{m_{i}}$ for each mass $m_{i}$, we are able to tackle with the AHW model for light unflavored mesons.
-III.1 AHW model for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
-We start applying the AHW model for mesons with even ${\cal CP}$, isospin $I=1$, and spins $S\geq 1$, taking the mass of the $\rho(770)$ state from PDG as an input, which fixes $\Lambda_{\rm QCD}=150.96$ MeV.
+The uncertainty $\delta m_{0}$ of the state with $S=1$ for each meson family in our model will be the computed as follows: we consider $\delta m_{0}$ in the family $\rho(770)$ as the arithmetic mean of all uncertainties in the PDG data for this family, which represents 1.3% of the mass of $\rho(770)$.
+In order to have a uniform criterium, we will take the uncertainties $\delta m_{0}$ of each family to be 1.3% of their $S=1$ state mass.
+$I^{G}(J^{PC})$ Meson $a_{\rm Log}$ AHW (Log) Regge Trajectory (GeV) $R^{2}$
+$1^{+}(1^{--})$ $\rho(770)$ $2.77\pm 0.12$ $J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03$ 0.9998
+$0^{-}(1^{--})$ $\omega(782)$ $2.61\pm 0.09$ $J=(0.90\pm 0.01)M^{2}+0.48\pm 0.03$ 0.9997
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $0.56\pm 0.03$ $J=(0.78\pm 0.02)M^{2}-0.13\pm 0.08$ 0.9985
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $0.95\pm 0.06$ $J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05$ 0.9993
+Table 4: The Regge trajectories in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the AHW model, Eqs.
+(20) and (21), with the corresponding values of $R^{2}$ from Eq.
+(24).
+The coefficients $a_{\rm Log}$ presented here come from the minimization of $Q$, Eq.
+(23).
+Now that we have defined our criteria for best fit and how we compute the uncertainties $\delta{m_{i}}$ for each calculated mass $m_{i}$, we are able to tackle the AHW model for light unflavored mesons.
+Figure 2: Regge trajectories ($J\times M^{2}$) for mesons, with masses $M$ expressed in GeV, from the AHW Log model (blue lines) using the minimization of $Q$, Eq.
+(23), and from PDG data (orange lines).
+Red dots are the values obtained from Eq.
+(21), while black dots from PDG data, both shown with uncertainties in Tables 3 and 1, respectively.
+The lines are obtained by linear regression of these points.
+Upper left panel: mesons with even ${PC}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
+We start applying the AHW model for mesons with even ${PC}$, isospin $I=1$, and spins $S\geq 1$, taking the mass of the $\rho(770)$ state from PDG as an input, which fixes $\Lambda_{\rm HW}=150.96$ MeV for this family.
 We minimize the parameter $Q$, given by Eq.
@@
 $\Delta_{\rm Log}(S)=(2.77\pm 0.12)\ln(S).$ (26)
-The masses with its uncertainties obtained within this model, from Eq.
-(23), with these anomalous dimensions, are presented in Table 3, together with percentual deviations in comparison with PDG data.
+The masses with their uncertainties obtained within this model, from Eq.
+(21), with these anomalous dimensions, are presented in Table 3, together with percentage deviations in comparison with PDG data.
 The value of $\Delta_{\rm Log}(S)$ of the above equation for each spin $S$ is also shown in the last column of this Table.
 The Regge trajectory obtained by doing linear regression with the obtained masses from this AHW Log model is given by
-$J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03,$ (28)
-with $\chi^{2}/ndf=4.32/4=1.08$.
+$J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03.$ (27)
 This trajectory is presented in the upper left panel of Fig.
-1 as the blue line, and can be compared with the Regge trajectory of this family from PDG data, shown in Table 2.
-The red dots are the values in the $J\times M^{2}$ plane of the mesons $\rho(770)$, $a_{2}(1320)$, $\rho_{3}(1690)$, $a_{4}(1970)$, $\rho_{5}(2350)$, and $a_{6}(2450)$ in this AHW Log model, with their uncertainties.
-III.2 AHW model for mesons with even ${\cal CP}$, $I=0$, and $S\geq 1$
-Here, we discuss the case of mesons with even ${\cal CP}$, isospin $I=0$, and spins $S\geq 1$.
-Minimizing $Q$, given by Eq.
-(25), we obtain $a=2.61\pm 0.09$ for the anomalous dimension, Eq.
-(22), which reads
-$\Delta_{\rm log}(S)=(2.61\pm 0.09)\ln(S).$ (29)
-Using the mass of the state $\omega(782)$ from PDG as an input, which implies $\Lambda_{\rm QCD}=152.24$ MeV for this case, we calculate the tower of masses with the uncertainties in this AHW Log model and show these results in Table 3.
-The percentual deviations of these masses in comparison with PDG are also presented, as well as the value of the anomalous dimension $\Delta_{\rm Log}$, for each meson state.
-The Regge trajectory obtained by doing linear regression with the masses obtained by the AHW Log model is given by
-$J=(0.90\pm 0.01)M^{2}+0.48\pm 0.03,$ (30)
-with $\chi^{2}/ndf=8.88/4=2.22$.
-This trajectory is presented in the upper right panel of Fig.
-1 as the blue line.
-The red dots are the values in the $J\times M^{2}$ plane of the mesons with even ${\cal CP}$, isospin $I=0$, and spins $S\geq 1$, in this AHW Log model, with their uncertainties.
-This trajectory can be compared with the one obtained from PDG for the $\omega(782)$ family, presented in Table 2.
-III.3 AHW model for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$
-Now we present our results for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$. The minimization of $Q$, Eq. (25), gives $a=0.56\pm 0.03$ for the the anomalous dimension, Eq. (22), which becomes
-$\Delta_{\rm Log}(S)=(0.56\pm 0.03)\ln(S).$ (31)
-The masses obtained in this AHW Log model are shown in Table 3, with their uncertainties, where we used the mass of the state $b_{1}(1235)$ as an input, so that $\Lambda_{\rm QCD}=239.41$ MeV.
-The Regge trajectory obtained by linear regression with these masses is given by
-$J=(0.78\pm 0.02)M^{2}-0.13\pm 0.08,$ (32)
-with $\chi^{2}$/ndf$=27/2=13.5$.
-This trajectory is represented in the lower left panel of Fig.
-1 as the blue line and can be compared with the one obtained from PDG data, shown in Table 2.
-The red dots are the values obtained from this AHW Log model, with their corresponding uncertainties.
-Table 3: Masses in MeV for mesons classified under even or odd $\mathcal{CP}$ with isospin $I=0,1$ and spins $S\geq 1$ from the AHW Log model, determined from Eq.
-(23), and the percentual deviations $\delta_{\rm Log}$ of the masses from this model compared with PDG data.
-For each family the mass of the state with $J=S=1$ is taken as an input from PDG.
-In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq.
+2 as the blue line, together with the one from PDG (orange line).
+The red dots are the values in the $J\times M^{2}$ plane of the mesons $\rho(770)$, $a_{2}(1320)$, $\rho_{3}(1690)$, $a_{4}(1970)$, $\rho_{5}(2350)$, and $a_{6}(2450)$ in this AHW Log model, with their uncertainties, while the black dots represent PDG data from Table 1.
+[…]
+For completeness and comparison, we also present in the Appendix VI a more standard minimization of $\chi^{2}$, defined in Eq.
 (33).
-$J^{PC}$ $\mathcal{CP}$ $I$ Meson AHW Log $\delta_{\rm Log}$ $\Delta_{\rm Log}$
-$1^{--}$ even 1 $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
-$2^{++}$ even 1 $a_{2}(1320)$ $1309.95\pm 22$ $0.53\%$ $1.92$
-$3^{--}$ even 1 $\rho_{3}(1690)$ $1681.02\pm 31$ $0.46\%$ $3.04$
-$4^{++}$ even 1 $a_{4}(1970)$ $1988.75\pm 37$ $1.11\%$ $3.84$
-$5^{--}$ even 1 $\rho_{5}(2350)$ $2262.37\pm 43$ $2.90\%$ $4.46$
-$6^{++}$ even 1 $a_{6}(2450)$ $2514.58\pm 47$ $2.64\%$ $4.96$
-$1^{--}$ even 0 $\omega(782)$ $782.66\pm 7.83$ $0\%$ 00
-$2^{++}$ even 0 $f_{2}(1270)$ $1302.58\pm 17$ $2.12\%$ $1.81$
-$3^{--}$ even 0 $\omega_{3}(1670)$ $1666.4\pm 23$ $0.04\%$ $2.87$
-$4^{++}$ even 0 $f_{4}(2050)$ $1969.63\pm 28$ $2.40\%$ $3.62$
-$5^{--}$ even 0 $\omega_{5}(2250)$ $2240.19\pm 32$ $0.44\%$ $4.20$
-$6^{++}$ even 0 $f_{6}(2510)$ $2490.23\pm 36$ $1.02\%$ $4.68$
-$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 5.0$ $0\%$ 00
-$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1640.57\pm 8.35$ $1.80\%$ $0.39$
-$3^{+-}$ odd 1 $b_{3}(2030)$ $1991.55\pm 11$ $1.99\%$ $0.62$
-$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2316.73\pm 13$ $2.97\%$ $0.78$
-$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 10$ $0\%$ 00
-$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1629.93\pm 18$ $0.80\%$ $0.66$
-$3^{+-}$ odd 0 $h_{3}(2025)$ $2003.12\pm 25$ $1.08\%$ $1.04$
-$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2339.01\pm 30$ $0.47\%$ $1.32$
-Figure 1: Regge trajectories (blue lines) for mesons, with masses $m$ expressed in GeV, from the AHW Log model.
-Red dots are the values obtained from Eq.
-(23) in the AHW Log model with uncertainties shown in Table 3.
-The blue lines are obtained by linear regression of these points.
-Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
-III.4 AHW model for mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$
-For mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$, we minimize the function $Q$, Eq. (25), finding $a=0.95\pm 0.06$ for the anomalous dimension, Eq. (22), obtaining
-$\Delta_{\rm Log}(S)=(0.95\pm 0.06)\ln(S).$ (33)
-The masses obtained in this AHW Log model are presented, with uncertainties, in Table 3, where we used the mass value of the state $h_{1}(1170)$ from PDG as an input, such that $\Lambda_{\rm QCD}=227.04$ MeV.
-The Regge trajectory obtained by doing linear regression with these masses is given by
-$J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05,$ (34)
-with $\chi^{2}$/ndf$=3.52/2=1.76.$ This trajectory can be compared with the one obtained from PDG data shown in Table 2 and is presented in the lower right panel of Fig.
-1 as the blue line.
-The red dots are the values of the masses obtained within this AHW Log model with their uncertainties.
+As one can check, the results of both minimizations are very close to each other (actually, they coincide for the $h_{1}(1170)$ family, within the precision we are working here).
 IV Linear AHW for mesons
-Recently, we have shown Costa-Silva:2023vuu that one can obtain asymptotically linear Regge trajectories for glueballs with even spins $S$ from an AHW model, when considering a total dimension $\Delta_{\rm Total}$ proportional to $\sqrt{S}$.
+Recently, we have shown Costa-Silva:2023vuu that one can obtain asymptotically linear Regge trajectories for glueballs with even spins $S$ from an AHW model, when considering a total dimension, $\Delta_{\rm Total}$, proportional to $\sqrt{S}$.
+By asymptotically linear Regge trajectories, we mean functional relations between $J$ and $M^{2}$ which are not exactly linear but can approximate very well the linear behavior with the desired precision.
 In order to find asymptotically linear Regge trajectories for light unflavored mesons with spins $S\geq 1$, we apply a similar procedure here with a small modification in the power of the spins $S$ as
-$\Delta_{\rm Lin}=a_{\rm Lin}(S^{\frac{1}{2}+\epsilon}-1)+4,$ (35)
+$\Delta_{\rm LAHW}=a_{\rm LAHW}(S^{\frac{1}{2}+\epsilon}-1)+4,$ (28)
 where in the ideal case, $\epsilon\rightarrow 0$.
-Note that when $S=1$, which are the lowest spin state for each mesonic family considered here, we recover the canonical dimension $\Delta=4$, compatible with Eq.
+Note that when $S=1$, which is the lowest spin state for each mesonic family considered here, we recover the canonical dimension $\Delta=4$, compatible with Eq.
 (9).
-Here, the parameter $a_{\rm Lin}$ is obtained exactly as in the AHW Log model, by minimizing $Q$ in Eq.
+$I^{G}(J^{PC})$ Meson Linear AHW $\delta_{\rm LAHW}$ $\Delta_{\rm LAHW}$
+$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
+$1^{-}(2^{++})$ $a_{2}(1320)$ $1296.75\pm 21$ $1.53\%$ $1.85$
+$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1678.61\pm 30$ $0.60\%$ $3.03$
+$1^{-}(4^{++})$ $a_{4}(1970)$ $1993.84\pm 38$ $1.36\%$ $3.87$
+$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2268.03\pm 44$ $2.66\%$ $4.49$
+$1^{-}(6^{++})$ $a_{6}(2450)$ $2513.71\pm 51$ $2.60\%$ $4.96$
+$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
+$0^{+}(2^{++})$ $f_{2}(1270)$ $1294.28\pm 20$ $1.47\%$ $1.76$
+$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1669.09\pm 29$ $0.13\%$ $2.88$
+$0^{+}(4^{++})$ $f_{4}(2050)$ $1978.53\pm 37$ $1.96\%$ $3.67$
+$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2247.67\pm 44$ $0.10\%$ $4.24$
+$0^{+}(6^{++})$ $f_{6}(2510)$ $2488.81\pm 50$ $0.97\%$ $4.67$
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
+$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1667.30\pm 26$ $0.20\%$ $0.48$
+$1^{+}(3^{+-})$ $b_{3}(2030)$ $2000.60\pm 37$ $1.55\%$ $0.65$
+$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2281.03\pm 46$ $1.38\%$ $0.65$
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
+$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1646.87\pm 26$ $1.85\%$ $0.72$
+$0^{-}(3^{+-})$ $h_{3}(2025)$ $2009.96\pm 37$ $0.74\%$ $1.07$
+$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2314.19\pm 47$ $0.59\%$ $1.22$
+Table 5: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the Linear AHW model, determined from Eq.
+(29), and the percentage deviations $\delta_{\rm LAHW}=\delta_{i}\times 100\%$, with $\delta_{i}$ given by Eq.
+(22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $Q$, Eq.
 (23).
-The parameter $\epsilon$ that represents a little deviation of the ideal case $\epsilon=0$ is chosen to obtain the best asymptotic linear Regge trajectory for each meson family.
+For each family the mass of the state with $J=S=1$ is taken as an input.
+In the last column we show the total dimension $\Delta_{\rm LAHW}$ for each state determined by Eq.
+(28).
+$I^{G}(J^{PC})$ Meson $a_{\rm LAHW}$ $\epsilon$ Linear AHW Regge Trajectory (GeV) $n$ $R^{2}$
+$1^{+}(1^{--})$ $\rho(770)$ $6.87\pm 0.16$ $0.00\pm 0.005$ $J=(0.87\pm 0.01)M^{2}+0.51\pm 0.02$ 6 0.999793
+$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01$ 50 0.999997
+$0^{-}(1^{--})$ $\omega(782)$ $6.67\pm 0.13$ $0.00\pm 0.005$ $J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02$ 6 0.999845
+$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01$ 50 0.999995
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $3.19\pm 0.03$ $0.05\pm 0.005$ $J=(0.81\pm 0.01)M^{2}-0.24\pm 0.03$ 4 0.999846
+$J=(0.8549\pm 0.0003)M^{2}-0.42\pm 0.01$ 50 0.999993
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $3.79\pm 0.08$ $0.04\pm 0.005$ $J=(0.75\pm 0.03)M^{2}-0.03\pm 0.01$ 4 0.999960
+$J=(0.7667\pm 0.0003)M^{2}-0.006\pm 0.01$ 50 0.999994
+Table 6: The Regge trajectories in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the Linear AHW model, Eq.
+(29).
+For each meson family we show two Regge trajectories corresponding to the linear regression of two sets of points with dimension $n$.
+These trajectories are plot in Figures 3 and 4.
+We also show the corresponding values of $R^{2}$ from Eq.
+(24).
+The coefficients $a_{\rm LAHW}$ and $\epsilon$ presented here come from the minimization of $Q$, Eq.
+(23).
+Here, the parameter $a_{\rm LAHW}$ is obtained exactly as in the AHW Log model, by minimizing $Q$ in Eq.
+(23).
+The parameter $\epsilon$ that represents a little deviation of the ideal case $\epsilon=0$ is chosen to obtain the best asymptotically linear Regge trajectory for each meson family.
 The value of $\epsilon$ is the one that minimizes the total deviation of the masses from the straight line representing the Regge trajectory.
 The masses that follows from this linear AHW model are determined by
-$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm QCD};\qquad\qquad(\nu=\Delta_{\rm Lin}-2),$ (36)
-where $\Delta_{\rm Lin}$ is defined in Eq.
+$m_{\nu,1}=\chi_{\nu,1}\Lambda_{\rm HW};\qquad\qquad(\nu=\Delta_{\rm LAHW}-2),$ (29)
+where $\Delta_{\rm LAHW}$ is defined in Eq.
 (28).
-Since $\Delta_{\rm Lin}=4$ for $S=1$, the values of $\Lambda_{\rm QCD}$ here in this model, for each meson family, are the same as those in the AHW Log model, as well as in the original HW presented in Table 2.
-As in the case of logarithmic anomalous dimension, we will analyze the families of light unflavored mesons with even or odd $\cal CP$, isospins $I=0,1$, and spins $S\geq 1$.
+Since $\Delta_{\rm LAHW}=4$ for $S=1$, the values of $\Lambda_{\rm HW}$ here in this model, for each meson family, are the same as those in the AHW Log model, as well as in the original HW presented in Table 2.
+As in the case of logarithmic anomalous dimension, we will analyze the families of light unflavored mesons with even or odd $PC$, isospins $I=0,1$, and spins $S\geq 1$.
 The Regge trajectories that we obtain within this model are asymptotically linear for each meson family with the appropriate choice of parameters.
-IV.1 Linear AHW for mesons with even ${\cal CP}$, $I=1$, and $S\geq 1$
-For mesons with even ${\cal CP}$, isospin $I=1$, and spin $S\geq 1$, the parameters in Eq. (35) that minimizes $Q$ are $a_{\rm Lin}=6.87\pm 0.16$ and $\epsilon=0$, so the total dimension reads
-$\Delta_{\rm Lin}=(6.87\pm 0.16)\sqrt{S}-2.87\pm 0.16.$ (37)
+Figure 3: Regge trajectories ($J\times M^{2}$) for mesons, with masses $M$ expressed in GeV, from the Linear AHW model (blue lines) with total dimension given by Eq.
+(28), using the minimization of $Q$, Eq.
+(23), and from PDG data (orange lines).
+Red dots are the values obtained from Eq.
+(29), while black dots come from PDG, both with uncertainties.
+The trajectories are obtained by linear regression of these points, for each meson family.
+Upper left panel: mesons with even $PC$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
+Figure 4: Regge trajectories ($J\times M^{2}$) for mesons, with masses $M$ expressed in GeV, from the linear AHW model, with total dimension given by Eq.
+(28), using the minimization of $Q$, Eq.
+(23).
+Red dots are the values obtained from Eq.
+(29) with uncertainties and the blue lines are the Regge trajectories, obtained by linear regression of these 50 points, for each meson family.
+Upper left panel: mesons with even ${PC}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${PC}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${PC}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${PC}$, $I=0$ ($h_{1}(1170)$ family).
+As in the previous section, we start with the $\rho(770)$ family, i.e., for mesons with even ${PC}$, isospin $I=1$, and spin $S\geq 1$, the parameters in Eq.
+(28) that minimize $Q$ are $a_{\rm LAHW}=6.87\pm 0.16$ and $\epsilon=0.00\pm 0.005$, so the total dimension reads
+$\Delta_{\rm LAHW}=(6.87\pm 0.16)\sqrt{S}-2.87\pm 0.16.$ (30)
 The masses obtained here, using Eq.
@@
 $J=(0.87\pm 0.01)M^{2}+0.51\pm 0.02,$ (31)
-with $\chi^{2}$/ndf$=8.68/4=2.17$ and is presented in the upper left panel of Fig. 2 as the blue line. The red dots represent these masses with their uncertainties.
-In order to show that the linear AHW model indeed present asymptotic linear Regge trajectories we also plot $J\times M^{2}$ for these mesons with even $\mathcal{CP}$, isospin $I=1$, and $S\geq 1$, for the first 50 states.
+and is presented in the upper left panel of Fig.
+3 as the blue line, together with the PDG data represented by black dots with Regge trajectory given by the orange line.
+The red dots represent the masses of the LAHW model with their uncertainties.
+In order to show that the linear AHW model indeed presents asymptotically linear Regge trajectories we also plot $J\times M^{2}$ for these mesons with even $PC$, isospin $I=1$, and $S\geq 1$, for the first 50 states with masses calculated from this model.
 The trajectory in this case becomes
-$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01,$ (39)
-with $\chi^{2}$/ndf$=72.96/48=1.52.$ This trajectory is presented in the upper left panel of Fig.
-3.
-Both trajectories can be compared with the one obtained from PDG data shown in Table 2 and are in good agreement with it.
-Table 4: Masses in MeV for mesons classified under even or odd $\mathcal{CP}$ with isospin $I=0,1$ and spins $S\geq 1$ from the Linear AHW model, determined from Eq.
-(36), and the percentual deviations $\delta_{\rm Lin}$ of the masses from this model compared with PDG data.
-For each family the mass of the state with $J=S=1$ is taken as an input.
-In the last column we show the total dimension $\Delta_{\rm Lin}$ for each state determined by Eq.
+$J=(0.8479\pm 0.0002)M^{2}+0.60\pm 0.01.$ (32)
+This trajectory is presented in the upper left panel of Fig.
+4.
+The trajectories from Eqs.
+(31) and (32) can be compared with the one obtained from PDG data shown in Table 1 and are in good agreement with it.
+These results are synthesized in Table 6.
+[…]
+For comparison, we also present in the Appendix VI the Tables 9 and 10, with the results of the minimization of $\chi^{2}$ applied to the LAHW model.
+As one can see from these tables, the results obtained from this method are very close to the minimization of $Q$, Eq.
+(23), within this model, shown in Tables 5 and 6.
+V Discusssion and Conclusions
+In this work, we proposed anomalous and linear HW models for light unflavored mesons.
+We have shown that the masses and Regge trajectories produced by these models are better than those obtained from the original HW and SW models, shown in Table 2 (see, e.g., Boschi-Filho:2002xih ; Boschi-Filho:2002wdj ; deTeramond:2005su ; Erlich:2005qh ; Boschi-Filho:2005xct ; Karch:2006pv ; Colangelo:2008us ).
+In fact, it is remarkable that the mass percentage deviations predicted by the AHW and LAHW models proposed here for mesons are less than $3\%$ when compared with PDG data.
+In particular, the linear AHW model produce asymptotically linear Regge trajectories, circumventing the well-kown problem of non-linearity of the original HW model.
+Since we analyzed four meson families classified under $PC$ parity (even or odd) and isospin $I$ values (0 or 1) starting with the states $\rho(770)$ (even $PC$, $I=1$), $\omega(782)$ (even $PC$ and $I=0$), $b_{1}(1235)$ (odd $PC$ and $I=1$), and $h_{1}(1170)$ (odd $PC$ and $I=0$) with the AHW and LAHW models it is possible to discuss some similarities and differences between these families.
+First of all, the similarities of the $\rho(770)$ and $\omega(782)$ families are very clear from the comparison of the parameters found by best fit to describe them.
+For instance, looking at Table 4 one sees that the coefficients $a_{\rm Log}$ for these two families are close to each other.
+The same can be said about the Regge trajectories equations appearing in this table.
+The same observation holds when analyzing Table 6, with respect to the parameters $a_{\rm LAHW}$ and $\epsilon$, and the corresponding trajectories.
+Since these two families share the same even $PC$ parity but different isospins, one can conclude that the dependence on the isospin of these parameters is small.
+The same conclusion is reached with the families of the $b_{1}(1235)$ and $h_{1}(1170)$ mesons which also have the same odd $PC$ parity, but different isospins.
+This discussion could also help us understand the role of the $\epsilon$ parameter in Eq.
 (28).
-$J^{PC}$ $\mathcal{CP}$ I Meson Linear AHW $\delta_{\rm LAHW}$ $\Delta_{\rm Lin}$
-$1^{--}$ even 1 $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
-$2^{++}$ even 1 $a_{2}(1320)$ $1296.75\pm 21$ $1.53\%$ $1.85$
-$3^{--}$ even 1 $\rho_{3}(1690)$ $1678.61\pm 30$ $0.60\%$ $3.03$
-$4^{++}$ even 1 $a_{4}(1970)$ $1993.84\pm 38$ $1.36\%$ $3.87$
-$5^{--}$ even 1 $\rho_{5}(2350)$ $2268.03\pm 44$ $2.66\%$ $4.49$
-$6^{++}$ even 1 $a_{6}(2450)$ $2513.71\pm 51$ $2.60\%$ $4.96$
-$1^{--}$ even 0 $\omega(782)$ $782.66\pm 7.83$ $0\%$ 00
-$2^{++}$ even 0 $f_{2}(1270)$ $1294.28\pm 16$ $1.47\%$ $1.76$
-$3^{--}$ even 0 $\omega_{3}(1670)$ $1669.09\pm 22$ $0.13\%$ $2.88$
-$4^{++}$ even 0 $f_{4}(2050)$ $1978.53\pm 28$ $1.96\%$ $3.67$
-$5^{--}$ even 0 $\omega_{5}(2250)$ $2247.67\pm 33$ $0.10\%$ $4.24$
-$6^{++}$ even 0 $f_{6}(2510)$ $2488.81\pm 38$ $0.97\%$ $4.67$
-$1^{+-}$ odd 1 $b_{1}(1235)$ $1229.5\pm 5.0$ $0\%$ 00
-$2^{-+}$ odd 1 $\pi_{2}(1670)$ $1667.30\pm 7.89$ $0.20\%$ $0.48$
-$3^{+-}$ odd 1 $b_{3}(2030)$ $2000.60\pm 10.75$ $1.55\%$ $0.65$
-$4^{-+}$ odd 1 $\pi_{4}(2250)$ $2281.03\pm 13.31$ $1.38\%$ $0.65$
-$1^{+-}$ odd 0 $h_{1}(1170)$ $1166\pm 10$ $0\%$ 00
-$2^{-+}$ odd 0 $\eta_{2}(1645)$ $1646.87\pm 17$ $1.85\%$ $0.72$
-$3^{+-}$ odd 0 $h_{3}(2025)$ $2009.96\pm 24$ $0.74\%$ $1.07$
-$4^{-+}$ odd 0 $\eta_{4}(2330)$ $2314.19\pm 31$ $0.59\%$ $1.22$
-Figure 2: Regge trajectories (blue lines) for mesons, with masses $m$ expressed in GeV, from the Linear AHW model, with total dimension given by Eq.
-(35).
-Red dots are the values obtained from Eq.
-(36) with uncertainties and the blue lines are obtained by linear regression of these points, for each meson family.
-Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
-Figure 3: Regge trajectories for mesons, with masses $m$ expressed in GeV, from the linear AHW model, with total dimension given by Eq.
-(35).
-Red dots are the values obtained from Eq.
-(36) with uncertainties and the blue lines are the Regge trajectories, obtained by linear regression of these 50 points, for each meson family.
-Upper left panel: mesons with even ${\cal CP}$, $I=1$ ($\rho(770)$ family); Upper right panel: mesons with even ${\cal CP}$, $I=0$ ($\omega(782)$ family); Lower left panel: mesons with odd ${\cal CP}$, $I=1$ ($b_{1}(1235)$ family); Lower right panel: mesons with odd ${\cal CP}$, $I=0$ ($h_{1}(1170)$ family).
-IV.2 Linear AHW for mesons with even ${\cal CP}$, $I=0$, and $S\geq 1$
-For mesons in the linear AHW model with even ${\cal CP}$, $I=1$, and $S\geq 1$, the parameters in Eq.
-(35) that minimizes $Q$ are $a_{\rm Lin}=6.67\pm 0.13$ and $\epsilon=0$.
-So, the total dimension reads
-$\Delta_{\rm Lin}=(6.67\pm 0.13)\sqrt{S}-2.67\pm 0.13.$ (40)
-The masses obtained by this model, using Eq.
-(36), are presented with respectives uncertainties in Table 4, taking the mass of the state $\omega(782)$ as an input from PDG.
-Furthermore, we also present in this table the total dimension and the percentual deviations from PDG data for each mesonic state.
-The Regge trajectory obtained by linear regression of these masses (first six states) is presented in the upper right panel of Fig.
-2 and is given by
-$J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02,$ (41)
-with $\chi^{2}$/ndf$=10.12/4=2.53.$ The upper right panel of Fig. 3 presents the Regge trajectory for the first 50 sates with even ${\cal CP}$, $I=0$, and $S\geq 1$, and it is given by
-$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01,$ (42)
-with $\chi^{2}$/ndf$=32.16/48=0.67.$ Both equations can be compared with the one obtained from PDG data, shown in Table 2, and are in good agreement with it.
-IV.3 Linear AHW for mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$
-For mesons with odd ${\cal CP}$, $I=1$, and $S\geq 1$ ($b_{1}(1235)$ family), the parameters in Eq.
-(35) that minimizes $Q$ are $a_{\rm Lin}=3.19\pm 0.03$ and $\epsilon=0.05$, so the total dimension becomes
-$\Delta_{\rm Lin}=(3.19\pm 0.03)S^{0.55}+0.81\pm 0.03.$ (43)
-The masses obtained in this model, using Eq. (36), are presented in Table 4. The Regge trajectory obtained for the first four points is presented in Fig. 2 and is given by
-$J=(0.81\pm 0.01)M^{2}-0.24\pm 0.03,$ (44)
-with $\chi^{2}$/ndf$=2.86/2=1.43.$
-For 50 points the trajectory is in Fig. 3 and is given by
-$J=(0.8549\pm 0.0003)M^{2}-0.42\pm 0.01,$ (45)
-with $\chi^{2}$/ndf$=163/48=3.39.$ Both equations, considering the uncertainties, are in agreement with the Regge trajectory obtained from PDG data, shown in Table 2.
-IV.4 Linear AHW for mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$
-The parameters in Eq.
-(35) that minimizes $Q$ for meson with mesons with odd ${\cal CP}$, $I=0$, and $S\geq 1$ ($h_{1}(1170)$ family) are $a_{\rm Lin}=3.79\pm 0.08$ and $\epsilon=0.04$, so the total dimension is
-$\Delta_{\rm Lin}=(3.79\pm 0.08)S^{0.54}+0.21\pm 0.08.$ (46)
-The masses obtained in this model, using Eq. (36), are presented in Table 4. The Regge trajectory for the first four points is presented in Fig. 2 and is
-$J=(0.7509\pm 0.033)M^{2}-0.03\pm 0.01,$ (47)
-with $\chi^{2}$/ndf$=0.22/2=0.11.$ The Fig. 3 presents the Regge trajectory for the first 50 points, given by
-$J=(0.7667\pm 0.0003)M^{2}-0.06\pm 0.01,$ (48)
-with $\chi^{2}$/ndf$=2.20/48=0.046.$ Both trajectories are in good agreement the one from PDG shown in Table 2.
-V Conclusions
-In this work, we proposed anomalous and linear HW models for light unflavored mesons.
-We have shown that the masses and Regge trajectories produced by these models are better than that obtained from the original HW and SW models.
-In particular, the linear AHW model produce asymptotically linear Regge trajectories, circumventing the well-kown problem of non-linearity of the original HW model.
-Finally, we expect to succesfully extend the analysis presented here to other particles as heavy and heavy-light mesons, as well as baryons. This is presently under investigation.
+Since we found $\epsilon=0$ for mesons with even $PC$ parity, while $\epsilon\sim 0.05$ in the case of odd $PC$ parity, it seems that this parameter is sensitive to these quantum numbers.
+This fact is related to the coefficients $\alpha^{\prime}$ and $\alpha_{0}$ of the linear Regge trajectories ($J=\alpha^{\prime}M^{2}+\alpha_{0}$) found for these mesonic families, in the linear anomalous HW model.
+In the case of even $PC$ parity, $\epsilon=0$, $\alpha^{\prime}\sim 0.9\,{\rm GeV}^{-2}$ and $\alpha_{0}\sim 0.50$, while for odd $PC$ parity one has $\epsilon\sim 0.05$, $\alpha^{\prime}\sim 0.8\,{\rm GeV}^{-2}$ and $\alpha_{0}<0$.
+From the physical point of view, one can say that the energy necessary to create a meson state with a given spin $S=J$  is lower for an even $PC$ parity one than the case with odd $PC$ parity with the same spin.
+This is in agreement with the experimental data from PDG, as can be seen in Table 1.
+Then, it is clear that the nonzero values of $\epsilon$ are related to an extra energy interaction required to form a mesonic state with odd $PC$ parity.
+In the case of the AHW Log model, an analogous discussion might be made looking at the values of the $a_{\rm Log}$ parameter, which is $\sim 2.7$ for the even parity case, while $\sim 0.75$ for the odd case.
+So one could speculate a relation like $a_{\rm Log}\sim 0.75+2\delta_{PC}$, where $\delta_{PC}$ is zero in the odd case (++-+ - or +-+- +) and 1 in the even case (+++++ + or --- -).
+A similar relation can also be proposed for the LAHW model with $a_{\rm LAHW}\sim 3.5+3.3\,\delta_{PC}=6.8-3.3\,\delta_{PC}$, which could be written more precisely in terms of the parameter $\epsilon$ as $a_{\rm LAHW}\sim 6.8-72\,\epsilon$.
+Finally, we expect to successfully extend the analysis presented here to other particles such as heavy and heavy-light mesons, as well as baryons. This is presently under investigation.
 Acknowledgments
@@
 is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under grant ##\## 310346/2023-1.
+VI Appendix: $\chi^{2}$ minimization
+$I^{G}(J^{PC})$ Meson AHW Log ($\chi^{2}$) $\delta_{\rm Log}$ $\Delta_{\rm Log}$
+$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
+$1^{-}(2^{++})$ $a_{2}(1320)$ $1308.72\pm 22$ $0.62\%$ $1.91$
+$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1679.12\pm 31$ $0.57\%$ $3.03$
+$1^{-}(4^{++})$ $a_{4}(1970)$ $1986.39\pm 37$ $0.98\%$ $3.83$
+$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2259.66\pm 43$ $3.02\%$ $4.44$
+$1^{-}(6^{++})$ $a_{6}(2450)$ $2511.59\pm 47$ $2.51\%$ $4.95$
+$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
+$0^{+}(2^{++})$ $f_{2}(1270)$ $1303.83\pm 22$ $2.22\%$ $1.82$
+$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1668.32\pm 30$ $0.08\%$ $2.88$
+$0^{+}(4^{++})$ $f_{4}(2050)$ $1972.01\pm 37$ $2.28\%$ $3.63$
+$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2242.93\pm 42$ $0.31\%$ $4.22$
+$0^{+}(6^{++})$ $f_{6}(2510)$ $2493.25\pm 46$ $1.15\%$ $4.69$
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
+$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1638.56\pm 27$ $1.92\%$ $0.38$
+$1^{+}(3^{+-})$ $b_{3}(2030)$ $1988.45\pm 37$ $2.14\%$ $0.60$
+$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2312.88\pm 44$ $2.79\%$ $0.76$
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
+$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1629.93\pm 25$ $0.80\%$ $0.66$
+$0^{-}(3^{+-})$ $h_{3}(2025)$ $2003.12\pm 33$ $1.08\%$ $1.04$
+$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2339.01\pm 39$ $0.47\%$ $1.32$
+Table 7: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the AHW Log model, determined from Eq.
+(21), and the percentage deviations $\delta_{\rm Log}=\delta_{i}\times 100\%$, with $\delta_{i}$ given by Eq.
+(22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $\chi^{2}$, Eq.
+(33).
+For each family the mass of the state with $J=S=1$ is taken as an input from PDG.
+In the last column we show the anomalous dimension $\Delta_{\rm Log}$ for each state determined by Eq.
+(20).
+We present here the more standard $\chi^{2}$ minimization as a complement to the $Q$ minimization discussed in section III, and also used in section IV to obtain the best fit of PDG data PDG:2022pth .
+As a second method to obtain the best fit of PDG data from the AHW models, we use the $\chi^{2}$ definition given by
+$\chi^{2}=\sum_{i}\frac{|m_{i}-M_{i}|^{2}}{M_{i}},$ (33)
+where $m_{i}$ is the mass obtained by our model and $M_{i}$ is the mass from PDG for a given state. So, according to this method, the best fit is such that $\chi^{2}$ given by (33) is minimum.
+$I^{G}(J^{PC})$ Meson $a_{\rm Log}$ AHW (Log) Regge Trajectory (GeV) $R^{2}$
+$1^{+}(1^{--})$ $\rho(770)$ $2.76\pm 0.12$ $J=(0.88\pm 0.01)M^{2}+0.50\pm 0.03$ 0.9997
+$0^{-}(1^{--})$ $\omega(782)$ $2.62\pm 0.11$ $J=(0.89\pm 0.01)M^{2}+0.48\pm 0.03$ 0.9997
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $0.55\pm 0.09$ $J=(0.78\pm 0.02)M^{2}-0.14\pm 0.08$ 0.9985
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $0.95\pm 0.07$ $J=(0.73\pm 0.01)M^{2}+0.04\pm 0.05$ 0.9993
+Table 8: The Regge trajectory in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the AHW model, Eqs.
+(20) and (21), using minimization with respect to the parameter $\chi^{2}$, Eq.
+(33).
+In the last column, we also show the value of $R^{2}$, defined in Eq.
+(24), for each trajectory.
+The results for the AHW model with logarithm anomalous dimensions, Section III, now minimized by the $\chi^{2}$ method are presented in Tables 7 and 8, and can be compared with the results obtained from the minimization of $Q$ shown in Tables 3 and 4.
+A comparison of these tables shows that the results of the two minimization methods are very close to each other for each family, and actually coincide for the $h_{1}(1170)$ family.
+Analogously, the minimization of $\chi^{2}$ can also be applied to the linear AHW model, discussed in Section IV.
+The results in this case are displayed in Tables 9 and 10.
+A comparison of these results with the ones from the $Q$ minimization, presented in Tables 5 and 6, shows that both methods produce similar outputs, an even exactly equal for the $\omega(782)$ family.
+$I^{G}(J^{PC})$ Meson Linear AHW ($\chi^{2}$) $\delta_{\rm LAHW}$ $\Delta_{\rm LAHW}$
+$1^{+}(1^{--})$ $\rho(770)$ $775.26\pm 10.08$ $0\%$ 00
+$1^{-}(2^{++})$ $a_{2}(1320)$ $1295.28\pm 20$ $1.64\%$ $1.84$
+$1^{+}(3^{--})$ $\rho_{3}(1690)$ $1676.08\pm 29$ $0.75\%$ $3.01$
+$1^{-}(4^{++})$ $a_{4}(1970)$ $1990.44\pm 37$ $1.19\%$ $3.85$
+$1^{+}(5^{--})$ $\rho_{5}(2350)$ $2263.87\pm 44$ $2.84\%$ $4.47$
+$1^{-}(6^{++})$ $a_{6}(2450)$ $2508.87\pm 50$ $2.40\%$ $4.93$
+$0^{-}(1^{--})$ $\omega(782)$ $782.66\pm 10.17$ $0\%$ 00
+$0^{+}(2^{++})$ $f_{2}(1270)$ $1294.28\pm 20$ $1.47\%$ $1.76$
+$0^{-}(3^{--})$ $\omega_{3}(1670)$ $1669.09\pm 29$ $0.13\%$ $2.88$
+$0^{+}(4^{++})$ $f_{4}(2050)$ $1978.53\pm 37$ $1.96\%$ $3.67$
+$0^{-}(5^{--})$ $\omega_{5}(2250)$ $2247.67\pm 44$ $0.10\%$ $4.24$
+$0^{+}(6^{++})$ $f_{6}(2510)$ $2488.81\pm 50$ $0.97\%$ $4.67$
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $1229.5\pm 16$ $0\%$ 00
+$1^{-}(2^{-+})$ $\pi_{2}(1670)$ $1665.95\pm 26$ $0.28\%$ $0.48$
+$1^{+}(3^{+-})$ $b_{3}(2030)$ $1998.25\pm 36$ $1.66\%$ $0.64$
+$1^{-}(4^{-+})$ $\pi_{4}(2250)$ $2277.85\pm 45$ $1.24\%$ $0.64$
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $1166\pm 15$ $0\%$ 00
+$0^{+}(2^{-+})$ $\eta_{2}(1645)$ $1649.35\pm 26$ $2.00\%$ $0.73$
+$0^{-}(3^{+-})$ $h_{3}(2025)$ $2014.27\pm 36$ $0.53\%$ $1.09$
+$0^{+}(4^{-+})$ $\eta_{4}(2330)$ $2320.03\pm 45$ $0.34\%$ $1.24$
+Table 9: Masses in MeV for mesons classified under even or odd $PC$ with isospins $I=0,1$ and spins $S\geq 1$, from the Linear AHW model, determined from Eq.
+(29), and the percentage deviations $\delta_{\rm LAHW}=\delta_{i}\times 100\%$, with $\delta_{i}$ from Eq.
+(22), of the masses from this model compared with PDG data, using minimization with respect to the parameter $\chi^{2}$, Eq.
+(33).
+For each family the mass of the state with $J=S=1$ is taken as an input.
+In the last column we show the total dimension $\Delta_{\rm LAHW}$ for each state determined by Eq.
+(28).
+$I^{G}(J^{PC})$ Meson $a_{\rm LAHW}$ $\epsilon$ Linear AHW Regge Trajectory (GeV) $n$ $R^{2}$
+$1^{+}(1^{--})$ $\rho(770)$ $6.85\pm 0.16$ $0.00\pm 0.005$ $J=(0.88\pm 0.01)M^{2}+0.51\pm 0.02$ 6 0.999799
+$J=(0.8524\pm 0.0002)M^{2}+0.60\pm 0.01$ 50 0.999997
+$0^{-}(1^{--})$ $\omega(782)$ $6.67\pm 0.13$ $0.00\pm 0.005$ $J=(0.89\pm 0.01)M^{2}+0.49\pm 0.02$ 6 0.999845
+$J=(0.8766\pm 0.0003)M^{2}+0.52\pm 0.01$ 50 0.999995
+$1^{+}(1^{+-})$ $b_{1}(1235)$ $3.18\pm 0.11$ $0.05\pm 0.005$ $J=(0.82\pm 0.01)M^{2}-0.25\pm 0.03$ 4 0.999841
+$J=(0.8595\pm 0.0003)M^{2}-0.44\pm 0.01$ 50 0.999993
+$0^{-}(1^{+-})$ $h_{1}(1170)$ $3.81\pm 0.12$ $0.04\pm 0.005$ $J=(0.75\pm 0.03)M^{2}-0.02\pm 0.01$ 4 0.999964
+$J=(0.7599\pm 0.0003)M^{2}-0.04\pm 0.01$ 50 0.999994
+Table 10: The Regge trajectories in GeV units for each light unflavored meson family (same $PC$ parity and same isospin $I$) from the Linear AHW model, Eq.
+(29), using the minimization with respect to the parameter $\chi^{2}$, Eq.
+(33).
+For each meson family we show two Regge trajectories corresponding to the linear regression of two sets of points with dimension $n$.
+In the last column, we also show the value of $R^{2}$, defined in Eq.
+(24), for each trajectory.